### v1

# Quantum field theory on curved manifolds

Abstract
Given a manifold of a system with internal degrees of freedom, such as Lorentz symmetry or gauge symmetry, the “curvature” is defined for the manifold. If one defines the local vacuum in the tangent space of the manifold, one can define a local mapping in the vicinity of the contact point, which is nothing but the Bogoliubov transformation. The curvature of the electromagnetic field gives the Schwinger effect, and the curvature of the base space introduces the local Unruh effect, which realizes Hawking radiation if applied on the Black hole horizon. To show how the Bogoliubov transformation appears on the manifold and why the local calculation is crucial there, we consider the Schwinger effect with a slowly varying electric field. Then, we show what happens in the Unruh effect if the acceleration is not a constant parameter.
1 Introduction
When the curvature of a manifold is given by a non-zero constant, a dynamical element of particle generation may be added to the system, even though the manifold itself seems to be static. The Schwinger effect[1] and Hawking radiation[2] are typical examples of such phenomena. Particle production in spacetime with curvature has been treated by various methods[3]. In this paper, we focus on the most primitive method for calculating Bogoliubov transformations: analysis with field equations. In this case, the Bogoliubov transformations are realized as the Stokes phenomena of the differential equation. In the study of particle production, when it is analyzed using field equations, it is customary to introduce asymptotic states (flat space at infinity in time or space) to define the “vacuum” and the particle number there. One might argue that the loss of locality is inevitable in such calculation, but at the same time, it is hard to believe that loss of locality is an essential feature of the phenomena. Indeed, if particle production is to be a heat bath, the loss of Markov property seems to be fatal for the analysis. In principle, it is also strange that the calculations cannot be completed at the place where the particles are created, without using the distant past or the distant future. Therefore, the loss of locality is clearly due to problems with the computational procedures. Turning to mathematics, gauge theory and general relativity are constructed by differential geometry and manifolds based on local analysis[4, 5, 6]. It is therefore natural to ask whether, using mathematical concepts based on local analysis, these problems can be solved without assuming asymptotic states at a distance, even if they are analyzed using field equations. In particular, in mathematical concepts of manifolds, tangent spaces appear naturally. These tangent spaces have properties quite similar to the asymptotic states. The reasons why a tangent space can be used as a vacuum are explained in detail in this paper. Since the tangent space is an essential feature of manifolds and differential geometry, and it has the properties required to define the vacuum in it, then it is very natural to assume that the vacuum has to be defined in the tangent space rather than for the asymptotic states. The aim of this paper is to rethink particle production on manifolds based on mathematical concepts. The best-known tangent space in physics is the local inertial system of general relativity. In the electromagnetic theory, it might be strange to define a tangent space, but indeed the tangent space of the principle bundle is commonly used in mathematics to define the curvature of the electromagnetic field. Our perspective is that particle production on manifolds, such as the Schwinger and the Unruh effects should be discussed in a unified way by using the same basic concepts of the differential geometry and manifolds. The crucial difference between the Unruh effect[7] and the Schwinger effect will be made clear in this paper.
When attempting to solve locally what has long been solved globally in physics, it is necessary to reconsider various mathematical properties that have previously been ignored. One might wonder why such mathematical concepts are required even in situations where solutions by path integrals are easily given. On this point, it would be important to emphasize here that path integral and Feynman diagrams are so cleverly constructed that a user can sometimes get the right result without having to worry about such mathematical concepts.
Section 2 of this paper describes the basic concepts of differential geometry and of manifolds, and the introduction of the symbols that are used in this paper. Section 3 describes the Schwinger effect when the electric field is constant or time-dependent. For fermions, it looks like the Landau-Zener transition[8]. By considering the case where the electric field is weakly depending on time, we show why the vacuum must be defined in the tangent space. This is the simplest example that clearly demonstrates that finding a naive global exact solution to an equation cannot always yield a physical solution. Section 4 discusses a local analysis of the Unruh effect and Hawking radiation. Again, as in the case of the Schwinger effect, the importance of defining the vacuum in the tangent space and the importance of local analysis is confirmed by considering the case of gradually varying acceleration.
2 An introduction to differential geometry and manifolds for particle production
First, we briefly describe the basic concepts of the manifold. A manifold is a fundamental concept in mathematics. It is defined as a topological space that locally resembles Euclidean space. This means that for every point in the manifold, there exists a neighborhood that can be mapped homeomorphically (i.e., through a continuous, bijective function with a continuous inverse) to an open subset of $\mathbb{R}^{n}$, where $n$ is the dimension of the manifold. For physics, this means that in principle the manifold will always have the required structure that is needed to define a local vacuum. [^2] To understand this more clearly for the Schwinger effect, we need to rethink tangent spaces after introducing gauge symmetries and matter fields. This local structure was designed from purely mathematical considerations. We will list here the most important definitions (and notations) of the manifolds that are convenient for our later discussions.
  1. 1.
    Local Euclidean Property: A manifold $M$ is an $n$-dimensional manifold if, for every point $p\in M$, there exists an open neighborhood $U_{i}$ of $p$ such that there is a homeomorphism $\varphi_{i}$ : $U_{i}\rightarrow V_{i}$, where $V_{i}$ is an open subset of $\mathbb{R}^{n}$. Note that $\bigcup_{i}U_{i}=M$. $\varphi_{i}(p)=(x^{1},x^{2},...,x^{n})$ is called a local coordinate of $p$ or simply a coordinate of $p$. $\varphi_{i}$ for $U_{i}$ is called the coordinate function of $U_{i}$.
  2. 2.
    Charts and Atlases:
    • A chart is a pair $(U_{i},\varphi_{i})$.
    • An atlas is a collection of charts that covers the entire manifold, allowing for transitions between different charts through coordinate transformations (using the coordinate functions).
Then, to describe the tangent space of a manifold, we define a tangent vector and a tangent space as follows:
  • A tangent vector: We introduce a tangent vector at $p$ as
    $X$ $=$ $\sum_{\mu}X^{\mu}\frac{\partial}{\partial x^{\mu}}.$ (2.1)
  • A tangent space: We introduce a tangent space as the space constructed by the whole tangent vectors at $p$. For concreteness, when the bases are given by $\{\frac{\partial}{\partial x^{1}},...,\frac{\partial}{\partial x^{n}}\}$, we have
    $T_{p}M$ $=$ $\left\{X^{\mu}\left.\frac{\partial}{\partial x^{\mu}}\right|X^{\mu}\in\mathbb{R}\right\},$ (2.2)
    where $T_{p}M\simeq\mathbb{R}^{n}$.
Then, a tangent bundle is defined by
$TM$$\coloneqq$ $\bigcup_{p\in M}T_{p}M.$ (2.3)
In this context, a vector field is a projection given by
$X\colon M$$\rightarrow$ $TM$
$p$$\rightarrow$ $X(p)\in T_{p}M.$ (2.4)
Also, the cotangent space at the point $p$, denoted as $T^{*}_{p}M$, is defined as the dual space of the tangent space $T_{p}M$. In this paper, both the tangent space and the cotangent space can simply be referred to as the tangent space if there is no possibility of confusion. For the sake of the explanation that follows, we are going to start with an intuitive explanation about defining the vacuum in the tangent space. We are now going to construct a field theory on $M$ as manifolds: the field theory constructed on $M$ is, of course, also defined at $p\in M$. Then, the theory defined on $p$ is naturally extended into the tangent space. A key feature of the theory extended to tangent spaces is that curvatures vanish. This is trivial with respect to the spacetime curvature, as the tangent space is obviously flat, but it should not be trivial with respect to gauge field curvature, as the gauge field curvature may not vanish in the flat spacetime. The vanishing curvature of the gauge field is understood by the local trivialisation that will be introduced in the definition of the fibre bundle in the following. Although a tangent space can be defined on any space defined as a manifold, here the vacuum is defined in the tangent space of $M$, as explained above.
Above, we have described tangent vector bundles to introduce the concept of bundles. It seems obvious that one can consider similar vector bundles in general (not necessarily for tangent vectors of $M$). This is the idea behind the fibre bundle. The coordinate functions defined in tangent vector bundles naturally had a direct product structure. [^3] Intuitively, the local direct product structure gives a point where one can start the discussion with a global symmetry (because they are given by the direct product) when describing a local symmetry. When considering general fibre bundles, the direct product structure has to be introduced by hand. To clarify the notation used, we refer to the definitions below.
  • Fibre bundle: A fibre bundle is defined by $(E,\pi,M,F,G)$, where
    1. 1.
      $E$ is the total space
    2. 2.
      $M$ is the base space
    3. 3.
      The projection $\pi$: $E\rightarrow M$ is a continuous surjection known as the projection map. The inverse $F\coloneqq\pi^{-1}(p)$ is called the fibre at $p\in M$.
    4. 4.
      The structure group $G$ acts on $F$ from the left.
    5. 5.
      The local trivialization is introduced by $\phi_{i}$ for the open coverings $\{U_{i}\}$ of $M$ as (sometimes $\phi_{i}^{-1}$ instead of $\phi_{i}$ is called the local trivialization because $\phi_{i}^{-1}$ seems to realize the direct product structure.)
      $\phi_{i}\colon U_{i}\times F$$\rightarrow$ $\pi^{-1}(U_{i})$
      $(p,f)$$\rightarrow$ $\phi_{i}(p,f),$ (2.5)
      where $E$ locally becomes a direct product.
Thanks to the introduction of a local direct product structure in the fibre, connections can be defined in a natural way. At this point, the introduction of a direct product structure may not seem to make sense as physics, but when one considers the role that local inertial systems actually play in general relativity, one has to consider that it does make sense as physics. Since in mathematics general relativity and gauge theories are described in a unified way, it seems to be natural to treat the two in a unified way. We believe that essential differences are highlighted only when analyses are held on the same foundations as far as possible.
Before introducing the connection, we should mention a scalar field used in the field theory of physics, which appears as a section of $E$ of a manifold. A “section” of a manifold will need some explanation. If we think of the elementary function $y=f(x)$ as a map $\mathbb{R}\rightarrow\mathbb{R}$, this function cuts $\mathbb{R}$ at the destination and gives a single value; if we draw a graph of $y=f(x)$ on the $xy$ space, its appearance will look like considering a section of the $xy$ space. Literally, a scalar field of the field theory corresponds to a section of the corresponding mapping. More abstractly, suppose that we are given some kind of mapping; If we give this mapping a concrete form, we are looking at a section. For cases where the mapping has internal degrees of freedom, a section is seen as one concrete form is selected. In physics, selecting the frame of a particular observer in special relativity is a section of the frame bundle.
It is particularly important not to confuse conventional “gauge fixing” with such a “section”. To understand the essence, recall the difference between the metric $g_{\mu\nu}$ in the theory of gravity and the gauge fixing of its fluctuation $\delta g_{\mu\nu}$. We will later analyze particle production in the vicinity of the contact point of a tangent space, but be careful not to confuse “the section used to define the contact point of the tangent space” with “gauge fixing of quantum fluctuations”.
Above, fibres have been introduced locally by means of local trivialization, and fibres on $U_{i}$ are laminated together to form a fibre bundle. When local trivialization varied from place to place, it was necessary to connect them by using the transformation. Therefore, it is very natural to ask whether the fibre bundle that is obtained in this simple way really has the correct differential structure. If one actually prepares a vector bundle [^4] The vector bundle has any dimension and is not always supposed to be the tangent vector bundle. and its section (e.g, a scalar field) and simply takes the derivative, it can be seen that the simple derivative is not covariant. Then, connections are introduced to solve this problem. This is called a covariant derivative. This connection is necessary for the differential geometry because fibre bundles are locally trivialized and then laminated together using transformations. If trivialization is possible in global, then inevitably the connections will disappear. [^5] On the other hand, if the section is defined for a moving frame, the connections do not vanish for the observer, even if the fibres are globally trivial on the manifold. The same is true for a “moving gauge” of the Schwinger effect. This point will be explained in more detail later. For the same reason, the value of the connection is supposed to vanish at the contact point of the tangent space.
To explain the situation using a concrete example, we consider a two-dimensional real vector field $\phi$: $M\rightarrow\mathbb{R}^{2}$ as a section of a vector bundle $E$, where $M=\mathbb{R}^{3}$ and $F=\mathbb{R}^{2}$. Since the fibre is a two-dimensional real space, we choose the structure group as $GL(2,\mathbb{R})$. Therefore, $\phi$ translates as $\phi^{\prime}(x)=g(x)\phi(x)$ by $g(x)\in GL(2,\mathbb{R})$ and laminated on the fibre. The question is if $d\phi(x)$ could also be a section of the bundle on which $GL(2,\mathbb{R})$ can act properly to be laminated on the fibre as the original function $\phi$ [^6] The exterior derivative is used here for simplicity of notation. The process of deriving the connection one-form ($A$) is the same as in field theory and is therefore omitted. What is important here is that by using
$\nabla\phi$ $=$ $(d+A)\phi,$ (2.6)
where
$A^{\prime}$ $=$ $-(dg)g^{-1}+gAg^{-1},$ (2.7)
one will see
$(\nabla\phi)^{\prime}$ $=$ $g\nabla\phi,$ (2.8)
which translates the same way as the original section (i.e, the scalar field $\phi$) of the vector bundle. A simple explanation is that since the two sections $\phi$ and $\nabla\phi$ transform in the same way, they can be laminated in the same way. This is the requirement for the consistency of the differential geometry. Here $A$ is called a connection or a gauge field. Although the transformation of the scalar field is trivial under the coordinate transformation, the effect of gravity is incorporated in a natural way as $d\phi$ is also a section of the frame bundle. The frame bundle is described below.
Given that Lorentz and gauge symmetries are treated as equivalence classes in field theory, one might think that the above discussion does not adequately describe the situation. To describe this point, it is necessary to introduce a principal bundle. The covariant derivative defined above can also be explained using the tangent space of the principal bundle. To give an overview without using further mathematical definitions, consider the simplest tangent space for an example. In the tangent space, there are various ways of taking coordinates depending on the coordinate transformation, so the principal bundle is the fibre that brings them all together. The principal bundle deals with such equivalence as a fibre where the coordinate transformation induces a motion on it. Note that in physics, the presence of an observer naturally defines a section of the principal bundle (frame bundle), since the observer chooses a unique frame. The importance of such a frame in physics (called a moving frame[9] for an accelerating observer) has already been confirmed by the Thomas precession[9] in a non-trivial way. In addition to the principal bundle, a spinor bundle is required if fermions are to be introduced[6]. However, further explanation is beyond the scope of this paper. The reader is referred to the relevant textbooks[4, 5, 6] for more details.
3 The Schwinger effect on manifolds
First, we consider the case where the curvature of the manifold is defined for a gauge symmetry. The simplest model uses the electromagnetic $U(1)$ gauge symmetry on a flat space-time and the curvature is introduced by a constant electric field. In this model, the manifold is static in the sense that the curvature is constant, but quantum theory expects a dynamical phenomenon on it. The particle production in this model is called the Schwinger effect[1]. In this case, the quickest way to avoid tedious discussions about manifolds is to use a powerful computational tool, the path integral[1, 10, 11]. However, in this paper, we venture a primitive analysis based on field equations to look closely at what happens on the manifold. The analysis on the manifold using a scalar field is already given in Ref.[12, 13, 14]. What is new in this paper is the analysis of the fermionic Schwinger effect as the Landau-Zener transition and its application to a time-dependent electric field. We show that the concept of the manifold is particularly important when considering a slowly varying electric field.
To understand the fermionic Schwinger effect as the Landau-Zener transition, we introduce the conventional decomposition of the Dirac fermion as[15, 16, 17]
$\psi$ $=$ $\int\frac{d^{3}k}{(2\pi)^{3}}e^{-i\bm{k}\cdot\bm{x}}\sum_{s}\left[u_{\bm{k},s}(t)a_{\bm{k},s}+v_{\bm{k},s}(t)b^{\dagger}_{-\bm{k},s}\right],$ (3.1)
where $v_{\bm{k},s}=C\left(\bar{u}_{\bm{k},s}\right)^{T}$, and $\psi$ obeys the single-field Dirac equation
$(i\hbar\not{\partial}-m)\psi$ $=$ $0$ (3.2)
Taking the momentum ${\bm{k}}\equiv k_{z}$ and defining [^7] Hereafter, we omit ${\bm{k}}$ in the indices of $u$.
$u_{s}$$\equiv$ $\left[\frac{u_{+}}{\sqrt{2}}\psi_{s},\frac{u_{-}}{\sqrt{2}}\psi_{s}\right]^{T},$
$v_{s}$$\equiv$ $\left[\frac{v_{+}}{\sqrt{2}}\psi_{s},\frac{v_{-}}{\sqrt{2}}\psi_{s}\right]^{T},$ (3.3)
where $\psi_{+}\equiv(1,0)^{T}$ and $\psi_{-}\equiv(0,1)^{T}$ are eigenvectors of the helicity operator. Carefully following the formalism given in Ref.[15], one will find
$\hbar\dot{u}_{\pm}$ $=$ $iku_{\mp}\mp imu_{\pm},$ (3.4)
which can be written in the matrix form as
$i\hbar\frac{d}{dt}\left(\begin{array}[]{c}u_{+}\\ u_{-}\end{array}\right)$ $=$ $\left(\begin{array}[]{cc}m&-k\\ -k&-m\end{array}\right)\left(\begin{array}[]{c}u_{+}\\ u_{-}\end{array}\right).$ (3.11)
Cosmological particle production after inflation has been discussed for time-dependent mass $m(t)$ in various situations. Such particle production is called preheating[18, 19]. It was first recognized in Ref.[20] that the Fermion preheating can be interpreted as the Landau-Zener transition[8], and the idea has been extended in Ref.[16, 17] to solve cosmological problems of particle production. Particle-antiparticle asymmetry is not discussed here, but when it is described by the multi-element Landau-Zener transition, there are seeds of asymmetry in the interference between different kinds of the Stokes phenomena[16, 17]. The relationship between cosmological particle production and the Landau-Zener transition is not discussed in detail in this paper, so more details and further explanations are left to these papers.
3.1 Constant electric field (constant curvature)
Introducing a constant electric field in the $z$-direction, we find
$\hbar\dot{u}_{\pm}$ $=$ $i(k+eE_{0}t)u_{\mp}\mp imu_{\pm},$ (3.12)
which can be written in the matrix form as
$i\hbar\frac{d}{dt}\left(\begin{array}[]{c}u_{+}\\ u_{-}\end{array}\right)$ $=$ $\left(\begin{array}[]{cc}m&-k-eE_{0}t\\ -k-eE_{0}t&-m\end{array}\right)\left(\begin{array}[]{c}u_{+}\\ u_{-}\end{array}\right).$ (3.19)
We will try to improve the analytical perspective by starting with a general formulation. We first consider the (generalized) Landau-Zener transition[8] with
$i\hbar\frac{d}{dt}\left(\begin{array}[]{c}X\\ Y\end{array}\right)$ $=$ $\left(\begin{array}[]{cc}D(t)&\Delta(t)^{*}\\ \Delta(t)&-D(t)\end{array}\right)\left(\begin{array}[]{c}X\\ Y\end{array}\right).$ (3.26)
Decoupling the equations, we have [^8] One might claim that the equation can be solved immediately using special functions. The reason for the somewhat roundabout approach here is that we want to examine the Stokes lines in the vicinity of the tangent space. To understand the structure of the Stokes lines, we use the Exact WKB (EWKB) developed in Refs.[21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
$\ddot{X}-\frac{\dot{\Delta}^{*}}{\Delta^{*}}\dot{X}+\left(-\frac{iD\dot{\Delta}^{*}}{\hbar\Delta^{*}}+\frac{i\dot{D}}{\hbar}+\frac{|\Delta|^{2}+D^{2}}{\hbar^{2}}\right)X=0.$ (3.27)
$\ddot{Y}-\frac{\dot{\Delta}}{\Delta}\dot{Y}+\left(\frac{iD\dot{\Delta}}{\hbar\Delta}-\frac{i\dot{D}}{\hbar}+\frac{|\Delta|^{2}+D^{2}}{\hbar^{2}}\right)Y=0.$ (3.28)
In the following only solutions of $X$ are examined. To obtain equations similar to the Schrodinger equation, we introduce $\hat{X}$ defined by
[…]
For the decoupled equations, the equation for $\hat{X}$ is
$\ddot{\hat{X}}+\left(\frac{-iD\dot{\Delta}^{*}}{\hbar\Delta^{*}}+\frac{i\dot{D}}{\hbar}+\frac{|\Delta|^{2}+D^{2}}{\hbar^{2}}+\frac{\ddot{\Delta}^{*}}{2\Delta^{*}}-\frac{3(\dot{\Delta}^{*})^{2}}{4(\Delta^{*})^{2}}\right)\hat{X}$ $=$ $0,$ (3.30)
which can be written as
$\hbar^{2}\ddot{\hat{X}}+\left(Q_{0}+\hbar Q_{1}+\hbar^{2}Q_{2}\right)\hat{X}$ $=$ $0,$ (3.31)
where
$Q_{0}$ $=$ $|\Delta|^{2}+D^{2}$
$Q_{1}$ $=$ $\frac{-iD\dot{\Delta}^{*}}{\Delta^{*}}+i\dot{D}$
$Q_{2}$ $=$ $\frac{\ddot{\Delta}^{*}}{2\Delta^{*}}-\frac{3(\dot{\Delta}^{*})^{2}}{4(\Delta^{*})^{2}}.$ (3.32)
Seeing the $\hbar$-dependence [^9] Our assumption here is that $\dot{\Delta}$ does not generate an additional factor of $\hbar$ or 1/$1/\hbar$ in the equation[20, 16]. , the Stokes lines of the above equation coincide with the simple equation[29, 31]
$\ddot{\hat{X}}$ $+$ $\frac{|\Delta|^{2}+D^{2}}{\hbar^{2}}\hat{X}=0.$ (3.33)
When $\Delta(t)=\lambda t$ and $D(t)=D_{0}$, the above equation is a well-known problem of quantum mechanics (i.e, scattering by an inverted quadratic potential).
There are a few things to be considered with care when using this equation for solving the Schwinger effect. The first and the most important is the definition of the vacuum. In the usual analysis of a constant electric field, one will find a scattering problem by an quadratic potential, where the two vacuum states (in and out states at $t=-\infty$ and $t=+\infty$ where the electric field is supposed to disappear) are defined as asymptotic states. Then, the Stokes phenomenon is assumed to occur between the two vacuum states. In this case, the gauge symmetry is used to explain the arbitrariness of positioning the top of the potential hill. However, as already explained, tangent spaces are naturally introduced in manifolds, and these tangent spaces have the property for defining the local vacuum states. Therefore, instead of losing the locality of the analysis by assuming the vacuum states at far away, we try to solve the problem by defining the vacuum on the tangent space. By definition, a feature of the tangent space is that the connection is zero in the space. This feature must be realized on the manifold at the point of contact. This is also a simple consequence of the local trivialization. This means that the gauge symmetry of each $U_{i}$ must be used carefully to make the connection vanish at the point. This makes the top of the quadratic potential coincide with the tangent space, as is illustrated in Fig.1.
Figure 1: The typical “potential” of the equation is shown for $V(t)=1-t^{2}$. The Stokes lines for the potential are shown on the complex $t$-plane. They are presented on the top. The Stokes lines appearing on $U_{i}$ of $M$ are illustrated in the bottom picture.
Then, they are laminated using gauge transformations. The situation is quite similar to the moving frame in special relativity. Similarly, since the tangent space is also a local inertial system, the velocity (or $k$) is assumed to be negligible to account for the inertial vacuum seen by the generated particle. [^10] In the Unruh effect, the vacuum is non-trivial when it is seen by the particle, not by someone else. Note that choosing a section of the observer (i.e, the inertial frame) does not kill the Lorentz symmetry of the theory. Similarly, the choice of the gauge in $U_{i}$ does not kill the gauge symmetry of the theory. This changes the definition of the electric field $E_{0}$ in the equation since the original $E_{0}$ was defined for an observer in the laboratory. Therefore, we denote the electric field in the equation described in the vicinity of the tangent space as $\hat{E}$ in the following calculations. Now we have defined everything in the vicinity of the contact point of the tangent space. Note that the “decoupling” of the equations performed above cannot be defined at the point of contact with the tangent space, as the non-diagonal element disappears from the matrix and the equations are already decoupled there. This shows that there is no mixing in the tangent space by definition. As discussed in the definition of manifolds, the Stokes phenomenon should be considered on an open set $U_{i}$ defined in a neighborhood of $p\in M$, and it is the physics around $p$ that is relevant for mixing solutions. Therefore, the Stokes phenomenon is considered here in a neighborhood where the point $\Delta=0$ is slightly avoided. The above equations show that the real-time axis traverses the Stokes line in the vicinity of the tangent space on every $U_{i}$[13]. See also Fig.1.
Thus, when the local setting is made natural as a manifold, it can be seen that the Stokes lines appear in the neighborhood of the local tangent space without ambiguity of the time and the gauge. This indicates that the Stokes Phenomenon is constantly mixing the vacuum solutions. Exactly speaking, the vacuum defined “in” the tangent space cannot describe the mixing by definition, while the mixing is seen in the neighborhood of the tangent space placed on each $U_{i}$. This is the same situation as the asymptotic states. In the same way as the correspondence with the vacuum is considered for the asymptotic states, the correspondence with the vacuum solution here is considered in the vicinity of the tangent space.
Here, one might notice that the frame prepared for the generated particles of the Schwinger effect actually represents an accelerated frame called the moving frame. For this simple reason, the analysis of the Schwinger effect would not be complete without an analysis of the Unruh effect. This topic will be discussed in the next section. In the remaining part of this section, we are going to look at the time-dependent case in a little more detail.
3.2 Time-dependent $E(t)$
We will now try to show an example where it makes no sense at all as physics to find a naive global solution to an equation defined on an open set $U_{i}$. It naturally depend on the situations whether a meaningful result can be obtained by such solutions when extrapolating the equations originally defined on $U_{i}$. However, almost all papers do not mention this point at all. In order to understand the problem without ambiguity, consider the case where the electric field changes gradually and there is no significant back-reaction from particle production. If the electric field oscillates, our assumption here is that the time period of the oscillation is much longer than the width of $U_{i}$. [^11] If the period of the oscillation is short, averaging over a period should be considered for the observation. In addition, the Stokes phenomenon in mathematics occurs on the Stokes “line” on the complex plane, but in physics, it will be better to assume an intrinsic width of the line. Therefore, the width of on an open set $U_{i}$ is assumed to be wider than the intrinsic width of the Stokes line.
Replacing $E_{z}(t)=E_{0}$ with $E_{z}(t)=\alpha t$, the problem becomes scattering by a “quartic” potential. The solution to this problem and the Stokes lines have been studied in great detail by Voros[22], and it has been found that the Stokes lines act away from the origin. See Fig.2 for more details about the Stokes lines. Details concerning conventional particle production can be found in Ref.[20]. Unlike the case of the constant electric field, the gauge ambiguity cannot naively shift the position of the quartic potential. Also, although intuitively the probability of particle production should change gradually if the electric field changes slowly, such calculation (scattering by the quartic potential solved for asymptotic states) will not change in such a way. Although it is immediately obvious to anyone who looks at the equation, we point out that the solutions of the equation, if they are used for the asymptotic states, can not solve the problem at all. It is easy to say that the setting of the problem is bad, but here we are going to try to clarify our understanding a little bit more. For our purpuses, it is important to go back to the definition of the manifold and solve the problem using local analysis.
Figure 2: The typical Stokes lines of the scattering problem by a quartic potential are shown for $V(t)=1-t^{4}$. In contrast to the quadratic potential, the Stokes lines are not crossing the origin.
Let us now reconsider the above question along the lines of the manifold and the differential geometry. First, consider an open set $U_{i}$ in the neighborhood of $t=t_{p}$ and let $A_{z}(t)=\frac{1}{2}\alpha t^{2}$ be expanded at $t=t_{p}$. Then, we have the “local potential” defined for $U_{i}$ as
$V(t)|_{U_{i}}$ $\simeq$ $e^{2}\hat{E}_{z}(t_{p})^{2}(t-t_{p})^{2},$ (3.34)
where by using the gauge of $U_{i}$ the connection is set to zero at the contact point $t=t_{p}$. This equation clearly meets the above requirement. In this way, the generation rate can be calculated for the local time ($t_{p}$) on the local open set ($U_{i}$).
There could be critical opinions about not defining asymptotic states for solving field equations to analyze particle production, but at least for cases like the one discussed here, the use of the tangent space seems to be inevitable. We would like to stress the importance of using two different definitions of the vacuum for different purposes. Furthermore, if the electric field changes much more rapidly and is treated as quantized, then the above treatment as a “background field” would no longer be appropriate and the field equations of the gauge field come into play. Then, we cannot imagine a better alternative to Feynman diagrams and path integrals.
As we have already seen in the simple example of a slowly changing electric field, it is very important and sometimes quite essential to solve the field equations locally on the manifold. Of course, when determining the averaged particle production rate in the case of repeated rapid oscillations or when the electric field is instantaneous, local evaluation of the particle production rate is useless for the experimental observation. Then, to get the required results, one should consider asymptotic states and averaging over a period. [^12] Analysis of these topics using the Stokes lines can be found in Ref.[33]. See also Refs.[34, 35].. What we have highlighted in this section is the case in which local calculation is needed to obtain correct results while a “naive” application of the conventional asymptotic states gives clearly a wrong answer. This distinction has never been explicitly recognized. We believe that the importance of the option of defining the vacuum in the tangent space has been made very clear by the simple model described in this section.
Let us now look at how such a definition of the vacuum has implications in relativity.
4 The Unruh effect and Hawking radiation on manifolds
The previous section dealt with the case where there is no qualitative ambiguity in defining the local vacuum, but in general relativity, even after the connection and the tangent spaces are determined, there are still further degrees of freedom left in the vierbein, which leaves ambiguity in defining the local vacuum. The most obvious difference is between the Lorenz frame and the local inertial frame. In mathematics, covariant derivatives are defined by using the Lorenz frame, which gives a vierbein that is diagonal with respect to the time-direction in the neighborhood. On the other hand, in physics, the vacuum is defined for the local inertial frame, which gives a vierbein that is diagonal with respect to the time-direction only at the contact point. In both cases, the tangent space is correctly defined at the point, but there is a difference in physics in the neighborhood.
To understand the situation more clearly, let us start by introducing the notion of the frame bundle in more detail. Suppose that $\varphi_{i}(p)$ is the coordinate function $\{x^{\mu}(p)\}$ of $U_{i}$. In the “coordinate basis”, $T_{p}M$ is spanned by $\{e_{\mu}\}=\{\partial/\partial x^{\mu}\}$, while the “non-coordinate bases” is explained as
$\hat{e}_{\alpha}$ $=$ $e_{\alpha}^{\mu}\frac{\partial}{\partial x^{\mu}},\,\,\,\,e_{\alpha}^{\mu}\in GL(m,\mathbb{R}),$ (4.1)
where the coefficients $e_{\alpha}^{\mu}$ are called vierbeins. Since $U_{i}$ is homeomorphic to an open subset $\varphi(U_{i})$ of $\mathbb{R}^{m}$ and each $T_{p}M$ is homeomorphic to $\mathbb{R}^{m}$, $TU_{i}\equiv\bigcup_{p\in U_{i}}T_{p}M$ is a $2m$-dimensional manifold, which can always be decomposed into a direct product $U_{i}\times\mathbb{R}^{m}$. This means that the local theory at that point (not in the neighborhood) is nothing but special relativity. Note that in differential geometry everything starts with local trivialization. Given a principal fibre bundle $P(M,G)$, one can define an associated fibre bundle as follows. [^13] The explanation here is in the opposite direction to the description of the principal bundle we have already given. Previously, we started from the fibre bundle to reach at the notion of the principal bundle. The explanation here is useful when the structure group is determined first. For $G$ acting on a manifold $F$ on the left, one can define an action of $g\in G$ on $P\times F$ by
$(u,f)$$\rightarrow$ $(ug,g^{-1}f)$ (4.2)
where $u\in P$ and $f\in F$. Now the associated fibre bundle is an equivalence class $P\times F/G$ in which $(u,f)$ and $(ug,g^{-1}f)$ are identified. For a point $u\in TU_{i}$, one can systematically decompose the information of $u$ into $p\in M$ and $V\in T_{p}M$. As we have mentioned, this leads to the projection $\pi$ : $TU_{i}\rightarrow U_{i}$. Normally, $\hat{e}_{\alpha}$ is requested to be orthonormal with respect to g;
$\mathrm{g}(\hat{e}_{\alpha},\hat{e}_{\beta})=e_{\alpha}^{\mu}e_{\beta}^{\nu}\mathrm{g}_{\mu\nu}=\delta_{\alpha\beta},$ (4.3)
where $\delta_{\alpha\beta}$ is replaced by $\eta_{\alpha\beta}$ for the Lorentzian manifold. The metric is obtained by reversing the equation
$\mathrm{g}_{\mu\nu}=e^{\alpha}_{\mu}e^{\beta}_{\nu}\delta_{\alpha\beta}.$ (4.4)
What is important for our discussion is that in an $m$-dimensional Riemannian manifold, the metric tensor $\mathrm{g}_{\mu\nu}$ has $m(m+1)/2$ degrees of freedom while the vielbein has $m^{2}$ degrees of freedom. For $m=4$, we have $10$ for the metric while $16$ for the vielbein. They are not identical. Each of the bases can be related to the other by the local orthogonal rotation $SO(m)$, while for the Lorentzian manifold it becomes $SO(m-1,1)$. The dimension of these Lie groups is given by the difference between the degrees of freedom of the vielbein and the metric. This shortly means that there are many (uncountable) choices for non-coordinate bases even after the metric is identified. This point will be very important when one looks at the Unruh effect[36, 37, 7]. The local inertial frame and the Lorentz frame have the same metric and are defining the same tangent space at the point. However, they are distinguished by the vierbein. The different choices of the “vacuum” is essential in the search for the Stokes phenomenon[13].
We describe the frame bundle further below. Associated with a tangent bundle $TM$ over $M$ is a principal bundle called the frame bundle $LM\equiv\bigcup_{p\in M}L_{p}M$ where $L_{p}M$ is the set of frames at $p$. The bundle $T_{p}M$ has a natural coordinate basis $\{\partial/\partial x^{\mu}\}$ on $U_{i}$ and a “frame” $u=\{X_{1},...,X_{m}\}$ at $p$ is expressed by the non-coordinate basis
$X_{\alpha}$ $=$ $X^{\mu}_{\alpha}\left.\frac{\partial}{\partial x^{\mu}}\right|_{p}$ (4.5)
where $(X^{\mu}_{\alpha})\in GL(m,\mathbb{R})$. If $\{X_{\alpha}\}$ is normalized by introducing a metric, the matrix $(X_{\alpha}^{\mu})$ becomes the vielbein.
The following point is very important for our discussion. A natural coordinate basis is prepared on the surface of $M$ and the inertial system is defined using a non-coordinate basis. This procedure naturally gives a notion of the “moving frame”[9], as the inertial frame seems to be rotated from time to time by the Lorentz transitions of the vielbeins, somewhat like spinning tea cups in amusement parks. Thomas precession is explained by the fact that multiple Lorentz transformations with different directions produce a rotation of the intrinsic space of the observer. If the observer stays in the same frame (no acceleration), there is no motion in the direction of the fibre of the frame bundle. In this case, the connection vanishes by definition because there is no transformation when laminating. Therefore, for an inertial observer, the distinction between coordinate and non-coordinate systems will be quite ambiguous. However, if one wants to describe an observer in accelerated motion on flat space-time, one has to define the local inertial frame for the observer, which moves on the frame bundle in a non-trivial way. This defines the section of the observer on the frame bundle. This means that on this section, one has to laminate the bundle by using non-trivial coordinate transformations. This (the observer’s non-trivial section of the frame bundle) introduces the connection for the observer, although the space-time is flat.
To be more specific, the vielbeins for constant acceleration ($a$) in the two-dimensional space-time at $\tau=\tau_{A}$ is
$(e_{A})_{\alpha}^{\mu}$ $=$ $\left(\begin{array}[]{cc}\cosh a(\tau-\tau_{A})&\sinh a(\tau-\tau_{A})\\ \sinh a(\tau-\tau_{A})&\cosh a(\tau-\tau_{A})\end{array}\right).$ (4.8)
Indeed, for such constant acceleration, the vielbeins have completely the same form for any time $(\tau)$. The transformation $(e_{A})_{\alpha}^{\mu}\rightarrow(e_{B})_{\alpha}^{\mu}$ on the frame bundle is the Lorentz transformation;
$L_{AB}$ $=$ $\left(\begin{array}[]{cc}\cosh a(\tau_{B}-\tau_{A})&-\sinh a(\tau_{B}-\tau_{A})\\ -\sinh a(\tau_{B}-\tau_{A})&\cosh a(\tau_{B}-\tau_{A})\end{array}\right).$ (4.11)
Note that the situation is very similar to the Schwinger effect for a constant electric field. In the Schwinger effect, the potential has been shifted by a gauge transformation; in the Unruh effect, the Lorentz transformation gives vielbeins of exactly the same shape. In both cases, the observer is always looking at the same physics on a static manifold.
Of course, the mathematical description of the frame bundle does not require an observer, but in physics an observer is inevitable. In this case, the observer is nothing but a section of the frame bundle. Previously, we have introduced $d\phi$ to describe covariant derivatives of the gauge symmetry. We have seen that both $\phi$ and $d\phi$ are the sections of the vector bundle, and the connection is introduced to make $d\phi$ covariant by the gauge symmetry. Now we can see that $d\phi$ is also a section of the frame bundle. If an observer stays in the same frame, there is no (non-trivial) connection required for $d\phi$. However, when an observer is accelerating, $d\phi$ traverses the frame bundle in the direction of the fibre and its motion causes Lorentz transformation. This is called “the moving frame”.
Is local analysis possible also for the Unruh effect, if we follow the previous calculations of the Schwinger effect? Our answer is “No”. What is important here is that the definition of the covariant derivative uses the Lorenz frame in which the vierbein is diagonalized in the neighborhood. On the other hand, as is shown explicitly above, the vierbein of the local inertial frame is diagonalized only at the point. (Note that the off-diagonal elements vanish since $\sinh a(\tau-\tau_{A})=0$ at $\tau=\tau_{A}$.) Since the covariant derivatives are defined for the Lorentz frame while the Unruh effect is defined for the inertial frame, it seems impossible to examine the Stokes phenomena of the Unruh effect directly (and locally) in terms of the field equations[13]. This mismatch has prevented the local analysis of the Unruh effect for a long time.
Now consider the physics that observers see in the Unruh effect. As far as the acceleration seen by the observer is constant, the physics seen by the observer is indistinguishable at any time due to the equivalence classes defined for the manifold. (More particularly, the observer feels the same vierbein all the time.) This is purely a mathematical consequence. If a dynamical effect (particle production) is manifested in such a situation, it must be explained by the vierbein of the local inertial frame. Extrapolating the coordinates to infinity and considering a global map as the Bogoliubov transformation is conceptually unacceptable, even if it could be plausible as a method[3]. In fact, it is known that such methods can lead to unnatural entanglements appearing between regions that should be uncorrelated. To argue the legitimacy of our local computation on manifolds, we must address this issue in this paper.
First of all, consider what happens if the scalar field equations were written down for an accelerating observer. Rindler coordinates are a specific set of coordinates used in the context of special relativity to describe the motion of an observer undergoing constant proper acceleration in flat spacetime. Thus, the Rindler coordinates form a coordinate chart that covers a specific region of Minkowski spacetime known as the Rindler wedge. Following the concept of manifolds, the simplest local inertial system is defined as a tangent space in the neighborhood of the point with zero velocity in Rindler’s coordinates. We start with the conventional Rindler metric[3, 9, 38];
$ds^{2}$ $=$ $-\left(1+ax\right)^{2}dt^{2}+dx^{2},$ (4.12)
where $x=a^{-1},t=0$ is the place for defining the simplest local inertial space. If one tries to define a local inertial system for any other point, one has to consider a distorted coordinate system due to Lorentz transformation. [^14] One can see more explanations in Ref.[9], in which figures for the moving frame can be found. As is immediately apparent from the metric, the field equation does not yield the local Stokes phenomenon as was the case with the Schwinger effect[12]. This indicates that there may be a fundamental error in the way of setting the problem. So far, many researchers have truncated locality here and moved on to find a solution in a global space[2, 3], but here we shall stick to the locality of the issue.
Since the metric (covariant derivatives) does not explain the local Stokes phenomenon, the only way left for us is to use the vierbein. The problem is that even if the vierbein is used for the calculation, the same trivial result is obtained once the field equations are written down. Therefore, the vierbein must be used without going through the field equations. Now consider what an accelerating observer would see if the observer looked directly at the vacuum (vacuum solutions) defined in the observer’s local inertial space. The vacuum is observed here by the particle itself, which is ejected from “the vacuum”. Since such particles have no momentum in their unique frame, we can neglect the $x$-dependent component of the vacuum solution. Considering $dt=\cosh\left(a\tau\right)d\tau$ from Eq.(4.8), we have for the time-dependent part;
[…]
which have the required Stokes phenomenon. See ref.[12, 14] if the reader is interested in the details about the Stokes phenomenon. One thing that should be noted is that the result does not meet the conventional (global) calculation by a factor of two. This point has to be explained.
To make the point clear, we consider the entanglements that appear in conventional calculations[3]. In our calculation, a difference of the factor of 2 appears in the non-perturbative factors emerging from the Stokes phenomenon. With entanglement, the probability ($P_{1}$ for a particle production) is squared because there are two particles produced. This gives $P_{entangled}=P_{1}^{2}$ in our calculation. In this way, our calculations will give the same results as the conventional calculations if the entanglement is included by hand. In short, if we “assume” entanglement, the two calculations coincide. So what happens if we use our calculations for Hawking radiation? Conventional calculations of the Unruh effect treat infinite inertial space as real, so the existence of such spaces on curved manifolds cannot be assumed. On the other hand, our calculations are based on the definition of manifolds, so the same calculations can be performed in the vicinity of the black hole horizon. When a pair of particles is created in the vicinity of the horizon, one particle inside the horizon can have negative energy (when it is seen from the outside), and the other in the outside can have positive energy. Again, as there are two particles produced, the total production probability is squared to obtain results consistent with Hawking radiation[12, 13]. Thus, if all calculations are performed faithfully to the definition of the manifold, strong doubts will arise about the entanglement of the Unruh effect. [^15] We are not claiming that our analysis is able to provide a proof that there is no entanglement in the global calculations. It is an indisputable fact that entanglement appears in the global calculation.
In the following, we will consider the Unruh effect when the acceleration changes, which is similar to the slowly changing electric field in the Schwinger effect. We are aware of papers claiming that the Unruh effect can be solved for various accelerating motions of the Unruh-DeWitt detector by adjusting the classical orbit of the detector. Again, our analysis cannot provide a proof that these results are wrong, but we believe we are giving a better alternative. Simply because the general calculations of the Unruh-DeWitt detector are somewhat different from the Unruh effect itself, we have treated them as different. See Ref.[14] for more details about what the local calculations of the Unruh-DeWitt detector look like if it is defined on a manifold, where we found the same factor of two discrepancy between the conventional global analysis and our local analysis. Note also that the Unruh-DeWitt detector requires an explicit interaction with the detector to be included, which is unlikely to be applicable directly to Hawking radiation.
4.1 When the acceleration rate is slowly time-dependent
For later calculations, let us first derive the Rindler coordinate for the case of constant acceleration using a somewhat roundabout approach. For simplicity, we restrict the motion to the x-axis direction. If the acceleration seen by the inertial system is $a(t)$ and the acceleration seen by an observer moving at the speed of $v(t)$ with respect to the inertial system is $a^{\prime}(t^{\prime})$, then the following relationship holds.
$a^{\prime}$ $=$ $\left(1-\frac{v^{2}}{c^{2}}\right)^{-\frac{3}{2}}a.$ (4.14)
Since the acceleration seen by an observer in accelerated motion with respect to an inertial system is $a^{\prime}$, $a^{\prime}$ is a constant in the conventional Unruh effect. After integrating both sides with respect to the time $t$, we find
$a_{0}t$ $=$ $\frac{v}{\left(1-\frac{v^{2}}{c^{2}}\right)^{\frac{1}{2}}},$ (4.15)
where $a(t)=dv/dt$ and $v(0)=0$ has been used. This ($v(0)=0$) means that the contact point with the tangent space is placed at $t=0$. Solving the above equation for $v(t)$, we find
$v(t)$ $=$ $\frac{a_{0}t}{\sqrt{1+\left(\frac{a_{0}t}{c}\right)^{2}}},$ (4.16)
which can be used to calculate the relation between the time coordinates as
[…]
$=$ $\frac{c}{a_{0}}\sinh^{-1}\left(\frac{a_{0}}{c}T\right).$
Finally, we find
$\frac{a_{0}}{c}T$ $=$ $\sinh\left(\frac{a_{0}}{c}T^{\prime}\right).$ (4.18)
Since the function $T$ of $T^{\prime}$ (i.e, $T(T^{\prime})$) is periodic in the direction of the imaginary axis of $T^{\prime}$, we can expect any vacuum function $F(T)$ described by the observer’s time $T^{\prime}$ to be periodic in the observer’s complex time. Intuitively, this suggests that what the accelerating observer sees in the vacuum is expected to be thermal. Here a simple question would arise. The periodic function for imaginary $T^{\prime}$ seen here is merely a parameterization of the elliptic function so that it takes an infinite limit for one of its double periodicity. The simple answer is that this may be a very special case due to the assumption that the acceleration is a constant. Let us see this point in more detail. In the above calculation, we simply had
[…]
where we set $C=0$ by $v(0)=0$. Let us relax the condition of this calculation and try to examine the following;
[…]
Then, for $f(t)=\alpha t+\beta t^{2}$ we have
[…]
which gives an elliptic integral after Legendre’s transformation and thus $T$ as the function of $T^{\prime}$ is described by an elliptic function. Obviously, if we consider $\beta\neq 0$, the periodicity of the function $T(T^{\prime})$ is not a simple imaginary. Such “coordinate systems” do not have the special properties of the Rindler coordinates. Therefore, Unlike the conventional Unruh effect for a constant acceleration, the vierbeins cannot be moved to the same form by the Lorentz transformation. This situation is quite similar to the case which appeared when we have considered the weakly time-dependent electric field in the Schwinger effect. Note that our local analysis only considers slices of the local elliptic function at the real axis and do not extrapolate it to infinity. The most familiar example of the elliptic function is probably the motion of a pendulum. Indeed, if we set
$f(t)$ $=$ $\int a_{0}\cos\omega t$ (4.22)
$=$ $\frac{a_{0}}{\omega}\sin\omega t\simeq a_{0}t$
for an “oscillation”, [^16] This “oscillation” is not defined for the observer’s time $t^{\prime}$. we can see that the approximate solution at $t=0$ is obtained for the Unruh effect in the similar way as a motion of a pendulum. [^17] Jacobi functions are complex-valued functions of a complex variable $z$ and a parameter $m(=k^{2})$. Using the elliptic integral of the first kind $K(m)$, the Jacobi $sn$ function has two periods $4K(m)$ and $2K(1-m)i$. In the present case we have $m<0$, while for an pendulum it becomes $m>0$. We comment on the case of solving it at other times. If we consider the Unruh effect at $t=t_{0}$, the inertial condition is now $v(t_{0})=0$. We thus have
$f(t)$ $=$ $\int a_{0}\cos\omega t$ (4.23)
$=$ $\frac{a_{0}}{\omega}\sin\omega t+C$
$\simeq$ $\left.\frac{a_{0}}{\omega}\sin\omega t\right|_{t=t_{0}}+\left.\left(\frac{a_{0}}{\omega}\sin\omega t\right)^{\prime}\right|_{t=t_{0}}(t-t_{0})+C,$
where the inertial condition gives $C=\frac{a_{0}}{\omega}\sin\omega t_{0}$. Finally, we have
$f(t)$ $\simeq$ $a^{\prime}(t_{0})(t-t_{0})$ (4.24)
for $a^{\prime}(t)=a\cos\omega t$ [^18] Here the primes are used in two different ways. The observer’s $a^{\prime}$ and $T^{\prime}$ should be distinguished from the derivatives. Using this result and the previous calculation for deriving $T(T^{\prime})$, one can find the Stokes phenomenon on the local space ($U_{i}$). These analyses explain how the periodicity of the Unruh effect in the imaginary $T^{\prime}$ direction is distorted by the time-dependent acceleration and how the Unruh effect can be calculated on a local space of the manifold without extrapolating the space to infinity.
5 Conclusions and Discussions
In this paper, the relationship between the Schwinger and Unruh effects has been discussed on the basis of their similarities as theories on manifolds. The two phenomena, which at first sight appear to be the same, turn out to be caused by completely different sources when one looks at the local structure of the manifold. As we expect negative reactions to the idea of defining the vacuum in the tangent space, this part of the article was explained in particular detail. In this paper, the case of a gradual change in curvature has been considered as an example where the correct answer can only be obtained when the vacuum is defined on the tangent space. In our local analysis, the entanglement of the Unruh effect appears to be an apparent one due to the extrapolation of the coordinate system, but this will require further multifaceted verification. The structure of the elliptic function for the complex time, manifested in the Unruh effect, is commonly found in curved spacetime. Further analysis is needed to determine what implication can be gained when such a structure is incorporated into a manifold.
In physics, it is well known that the connection of the electromagnetic $U(1)$ and the metric can be literally unified into the Kaluza-Klein theory. Although in our discussion of the Unruh effect it might have seemed that the Stokes phenomenon was not able to be derived from the metric, it is clear from the Kaluza-Klein theory that it can appear from the metric once they are actually embedded in the Kaluza-Klein theory. On the other hand, concerning the Unruh effect, the situation remains the same in Kaluza-Klein theory, as the Unruh effect is still hidden in analyses using covariant derivatives. The underlying cause of the Unruh effect is that the inertial system is attached in a twisted manner, which cannot be detected by the covariant derivatives (at least locally).
We hope that the local analyses presented in this paper will help people understand the physics of the quantum field theory on curved manifolds.

### v5

@@
 Abstract
-Given a manifold of a system with internal degrees of freedom, such as Lorentz symmetry or gauge symmetry, the “curvature” is defined for the manifold.
-If one defines the local vacuum in the tangent space of the manifold, one can define a local mapping in the vicinity of the contact point, which is nothing but the Bogoliubov transformation.
-The curvature of the electromagnetic field gives the Schwinger effect, and the curvature of the base space introduces the local Unruh effect, which realizes Hawking radiation if applied on the Black hole horizon.
-To show how the Bogoliubov transformation appears on the manifold and why the local calculation is crucial there, we consider the Schwinger effect with a slowly varying electric field.
-Then, we show what happens in the Unruh effect if the acceleration is not a constant parameter.
+This paper discusses how particle production from the vacuum can be explained by local analysis when the field theory is defined by differential geometry on curved manifolds.
+We have performed the local analysis in a mathematically rigorous way, respecting the Markov property.
+The exact WKB is used as a tool for extracting non-perturbative effect from the local system.
+After a serious application of the differential geometry and the exact WKB to particle production, we show that entanglement does not appear in the Unruh effect as far as the standard formulation by the differential geometry is valid.
+This result should not be attributed to a consistency problem between the “entanglement state” and the “standard field theory by differential geometry”, but to the fact that the conventional calculation of the Unruh effect is done by extrapolation which is not consistent with the differential geometry.
+The situation is similar to that of the Dirac monopole, but topology is not relevant and the basis for building field theories in differential geometry is strongly involved.
 1 Introduction
+The knowledge gained by constructing field theory by differential geometry is significant.
+The most famous is what happned in 70’s to the solution of the Dirac monopole[1, 2], where the differential geometry revealed that at least two solutions are involved for the Dirac monopole.
+The solution is famous in topology, but the mathematician’s deep understanding of local analysis in differential geometry is at the root of this story.
+Similar discoveries are likely to continue in the future, as differential geometry can involve various mathematical concepts to construct the field theory in a variety of sophisticated ways.
+In this paper, the Schwinger effect[3] and the Unruh effect[5, 4, 6] are analysed using the standard formulation of the field theory in terms of differential geometry.
+We paid particular attention to the Markov property and taken great care to ensure that the computation is localised and consistent with differential geometry.
 When the curvature of a manifold is given by a non-zero constant, a dynamical element of particle generation may be added to the system, even though the manifold itself seems to be static.
 The Schwinger effect[3] and Hawking radiation[4] are typical examples of such phenomena.
-Particle production in spacetime with curvature has been treated by various methods[3].
+Particle production in spacetime with curvature has been treated by various methods[6], but we are dissatisfied with conventional calculations because of loss of locality.
 In this paper, we focus on the most primitive method for calculating Bogoliubov transformations: analysis with field equations.
 In this case, the Bogoliubov transformations are realized as the Stokes phenomena of the differential equation.
+The main reason for us to stick to differential equations is that they are compatible with differential geometry.
 In the study of particle production, when it is analyzed using field equations, it is customary to introduce asymptotic states (flat space at infinity in time or space) to define the “vacuum” and the particle number there.
@@
 In principle, it is also strange that the calculations cannot be completed at the place where the particles are created, without using the distant past or the distant future.
-Therefore, the loss of locality is clearly due to problems with the computational procedures.
-Turning to mathematics, gauge theory and general relativity are constructed by differential geometry and manifolds based on local analysis[4, 5, 6].
-It is therefore natural to ask whether, using mathematical concepts based on local analysis, these problems can be solved without assuming asymptotic states at a distance, even if they are analyzed using field equations.
+Therefore, the loss of locality is clearly due to “for convenience”, which is the problem with the computational procedures of the physics side.
+Turning to mathematics, gauge theory and general relativity are constructed by differential geometry and manifolds, where local consistency is quite rigorous[7, 8, 9].
+It is therefore natural to ask whether, using mathematical concepts, these problems can be solved without assuming asymptotic states at a distance, or without introducing extrapolation.
+It could be difficult to notice the unnaturalness we are discussing about when one is accustomed to ordinary calculations in physics.
+However, we strongly insist that loss of locality is a serious problem that should be resolved by mathematics.
 In particular, in mathematical concepts of manifolds, tangent spaces appear naturally.
-These tangent spaces have properties quite similar to the asymptotic states.
-The reasons why a tangent space can be used as a vacuum are explained in detail in this paper.
-Since the tangent space is an essential feature of manifolds and differential geometry, and it has the properties required to define the vacuum in it, then it is very natural to assume that the vacuum has to be defined in the tangent space rather than for the asymptotic states.
-The aim of this paper is to rethink particle production on manifolds based on mathematical concepts.
-The best-known tangent space in physics is the local inertial system of general relativity.
+These tangent spaces have properties quite similar to the asymptotic states but appear in local.
+To explain in more detail, an operation called local trivialization is required at the point of contact with the tangent space, and this operation ensures that the connection is always zero at that point.
+The most familiar tangent space in physics will be the local inertial system, which is of course important in physics.222 To be more precise, the “inertial system” mentioned above is the tangent space with the observer’s (subjective) inertial frame.
+(The “frame” in the differential geometry is defined on the tangent space.) In physics, it is known that the subjective inertial frame causes Thomas precession[12].
+The reasons why a tangent space can be used as a vacuum are explained in detail in this paper, considering a simple example where the vacuum must be defined in the local tangent space.
+Indeed, considering how to introduce Lorentz symmetry into the local system, it is obvious that a vacuum must be defined in the tangent space even if it is defined as an asymptotic state at a distance.
+This point will be clarified in the discussion of charts and frames.
+Since both Lorentz and gauge symmetries are treated equivalently in the framework of differential geometry, it is quite natural to assume that the tangent space to the principal bundle of a gauge transformation has the same physical meaning.
+Note that the particle production discussed here is not an “overtly dynamical phenomenon” like prereheating after inflation, but rather a “phenomenon that causes dynamical particle production despite its apparent static state”.
+We will explain later that for the latter case a non-trivial treatment regarding gauge and Lorentz symmetries is required, which is difficult to be described without the help of differential geometry.
+In this paper, we will try to explain mathematical ideas as intuitive as possible, but since the concepts of the mathematics are rather complex, at least a basic knowledge of a typical differential geometry course is required to understanding the calculations correctly.
+Unlike differential geometry, which could be familiar as an undergraduate mathematics course, it would be very difficult to understand the exact WKB without reading any of the references.
+The serious difficulty with the exact WKB is that it looks so much like a regular WKB approximation that it is sometimes very hard to imagine how different the mathematics behind it is.
+The aim of this paper is to rethink locality of particle production, keeping the discussions as faithful as possible to mathematics.
+The main obstacle to the local description of non-perturbative particle production has been to find a local analysis of non-perturbative effect, which is now solved by the exact WKB in mathematics.
+The best-known tangent space in physics would be the local inertial system of general relativity.
 In the electromagnetic theory, it might be strange to define a tangent space, but indeed the tangent space of the principle bundle is commonly used in mathematics to define the curvature of the electromagnetic field.
 Our perspective is that particle production on manifolds, such as the Schwinger and the Unruh effects should be discussed in a unified way by using the same basic concepts of the differential geometry and manifolds.
-The crucial difference between the Unruh effect[7] and the Schwinger effect will be made clear in this paper.
+The crucial difference between the Unruh effect[5] and the Schwinger effect will be made clear in this paper, by using local calculations faithful to mathematics.
 When attempting to solve locally what has long been solved globally in physics, it is necessary to reconsider various mathematical properties that have previously been ignored.
@@
 On this point, it would be important to emphasize here that path integral and Feynman diagrams are so cleverly constructed that a user can sometimes get the right result without having to worry about such mathematical concepts.
+If the computation by path integrals is locally complete, our analysis will not lead to any new conclusions.
 Section 2 of this paper describes the basic concepts of differential geometry and of manifolds, and the introduction of the symbols that are used in this paper.
+Since differential geometry has almost never been seriously discussed in the analysis of the Unruh effect, we will be particularly careful to explain which aspects we will focus on.
+As the concept of manifolds is not explained from scratch, the reader who is not familiar with the mathematical approach is advised to read the references.
 Section 3 describes the Schwinger effect when the electric field is constant or time-dependent.
 For fermions, it looks like the Landau-Zener transition[11].
-By considering the case where the electric field is weakly depending on time, we show why the vacuum must be defined in the tangent space.
-This is the simplest example that clearly demonstrates that finding a naive global exact solution to an equation cannot always yield a physical solution.
+By considering the case where the electric field is weakly depending on time, we show why the vacuum must be defined in the local tangent space.
+This is the simplest example that clearly demonstrates that finding a naive global exact solution to an equation cannot always be a correct answer.
+Note that the basic concept explaining that a single solution cannot cover the whole space is the same as that of the Dirac monopole, although it has nothing to do with topology.
+In order to understand the causes of this non-trivial situation, it is necessary to take a closer look at the treatment of gauge and Lorentz symmetries in differential geometry, especially at the necessity of local trivialization.
+At the time of the Dirac monopole, much attention was paid to topology, but in the root of the discussions there has been the basic procedures in finding local solutions in differential geometry.
 Section 4 discusses a local analysis of the Unruh effect and Hawking radiation.
-Again, as in the case of the Schwinger effect, the importance of defining the vacuum in the tangent space and the importance of local analysis is confirmed by considering the case of gradually varying acceleration.
+Unlike the Unruh effect, where entangled pair production is expected at distant wedges, Hawking radiation is a pair production localised at the horizon.
+Therefore, we can calculate local particle production of Hawking radiation rigorously using differential geometry and the exact WKB, without indication of any discrepancy between the standard result.
 2 An introduction to differential geometry and manifolds for particle production
-First, we briefly describe the basic concepts of the manifold.
+First, we briefly describe the basic concepts of differential geometry and manifolds for the field theory.
+For our argument in this paper, a clear distinction between “charts” and “frames” is most important.
+Therefore, the main purpose of this introduction is to explain clearly why they are distinguished in mathematics and what are the implications for physics.
+We know that in physics the vacuum is exactly the same for observers in any frame.
+In mathematics, the Lie algebra and bundles are used to define the situation.
+However, the Lie algebra and the bundles are not defined for “charts”.
+As the Rindler coordinate system is a “chart”, it was (at least in principle) unnatural to define a “vacuum” in the Rindler chart to discuss the Bogoliubov transformation.
+Furthermore, the “moving frame” of an observer has its applicable range.
+The range is estimated as $\propto 1/a$ in Ref.[12], where $a$ is the rate of acceleration.
+The extrapolation beyond this range is not rigorous, but such extrapolation is normally used in the conventional calculation of the Unruh effect.
+Therefore, there is a need for a method that allows rigorous calculations to be made in local.
+Our idea is to introduce the exact WKB for the local analysis, using the vierbeins.
+Without understanding the unnaturalness described here, it will be hard to understand the following arguments.
+Although we will try to provide an overview as intuitive as possible, a basic knowledge of ordinary differential geometry is essential for a proper understanding of the subject.
+The standard description used here can be found in Ref.[7, 8, 9], but in principle different (and more complicated) extensions of manifolds could be possible in the description of the field theory[10].
+The mathematical concepts associated with field theory are not always confined to the basic scenario presented here.
 A manifold is a fundamental concept in mathematics.
@@
 This means that for every point in the manifold, there exists a neighborhood that can be mapped homeomorphically (i.e., through a continuous, bijective function with a continuous inverse) to an open subset of $\mathbb{R}^{n}$, where $n$ is the dimension of the manifold.
-For physics, this means that in principle the manifold will always have the required structure that is needed to define a local vacuum.
+For physics, this implies that in principle the manifold will always have the required structure that is needed to define a local vacuum.
 [^3] To understand this more clearly for the Schwinger effect, we need to rethink tangent spaces after introducing gauge symmetries and matter fields.
-This local structure was designed from purely mathematical considerations.
+The local structure of the differential geometry was designed from purely mathematical considerations.
 We will list here the most important definitions (and notations) of the manifolds that are convenient for our later discussions.
+In field theory, it is not possible to explain everything with just one manifold.
+Such complex structure is important for introducing gauge and Lorentz symmetries.
   1. 1.
@@
     Charts and Atlases:
-    • A chart is a pair $(U_{i},\varphi_{i})$.
+    • A chart is a pair $(U_{i},\varphi_{i})$. (See fig.1.)
     • An atlas is a collection of charts that covers the entire manifold, allowing for transitions between different charts through coordinate transformations (using the coordinate functions).
+Figure 1: A local coordinate function $\varphi_{i}$ is shown as a homeomorphism from $U_{i}\in M$ to $V_{i}\in\mathbb{R}^{n}$.
+A chart is a pair $(U_{i},\varphi_{1})$.
+The Rindler coordinate is a typical chart on flat spacetime.
+The “chart” must be discriminated from the “frame”.
 Then, to describe the tangent space of a manifold, we define a tangent vector and a tangent space as follows:
@@
     where $T_{p}M\simeq\mathbb{R}^{n}$.
-Then, a tangent bundle is defined by
+Then, a tangent bundle is defined by [^4] The tangent bundle is also a $2n$-dimensional manifold.
 $TM$$\coloneqq$ $\bigcup_{p\in M}T_{p}M.$ (2.3)
@@
 $p$$\rightarrow$ $X(p)\in T_{p}M.$ (2.4)
+Suppose that $\varphi_{i}(p)$ is the coordinate function $\{x^{\mu}(p)\}$ of $U_{i}$.
+In the “coordinate basis”, $T_{p}M$ is spanned by $\{e_{\mu}\}=\{\partial/\partial x^{\mu}\}$, while the “non-coordinate bases” is explained as
+$\hat{e}_{\alpha}$ $=$ $e_{\alpha}^{\mu}\frac{\partial}{\partial x^{\mu}},\,\,\,\,e_{\alpha}^{\mu}\in GL(m,\mathbb{R}),$ (2.5)
+where the coefficients $e_{\alpha}^{\mu}$ are called vierbeins.
+Note that the Lie brackets can be introduced only for the non-coordinate bases where the vierbeins play an essential role, as is shown in Fig.2.
+The figure shows that the vacuum must be defined in the tangent space, even if it is defined as an asymptotic state.
+Figure 2: The tangent space $T_{p}M$ is spanned by $\{e_{\mu}\}=\{\partial/\partial x^{\mu}\}$, where $\{x^{\mu}\}$ is the coordinate function of $U_{i}$.
+As is shown in Eq.(2.5) and the above picture, the vierbeins are essential for defining the Lie algebra.
+This means that the vacuum must be defined in the tangent space as far as the vacuum respects the Lorentz symmetry, even if it is placed at far distance as to describe the asymptotic state.
+The frame of an accelerating observer is called a moving frame, as it (the vierbeins) is changed by the Lorentz transformation time to time.
 Also, the cotangent space at the point $p$, denoted as $T^{*}_{p}M$, is defined as the dual space of the tangent space $T_{p}M$.
@@
 For the sake of the explanation that follows, we are going to start with an intuitive explanation about defining the vacuum in the tangent space.
-We are now going to construct a field theory on $M$ as manifolds: the field theory constructed on $M$ is, of course, also defined at $p\in M$.
+We are now going to construct a field theory on $M$ as manifolds [^5] We need many kinds of manifolds at the same time to explain the basic properties of the field theory.: the field theory constructed on $M$ is, of course, also defined at $p\in M$.
 Then, the theory defined on $p$ is naturally extended into the tangent space.
-A key feature of the theory extended to tangent spaces is that curvatures vanish.
+A key feature of the theory extended in the tangent space is that curvatures vanish.
 This is trivial with respect to the spacetime curvature, as the tangent space is obviously flat, but it should not be trivial with respect to gauge field curvature, as the gauge field curvature may not vanish in the flat spacetime.
-The vanishing curvature of the gauge field is understood by the local trivialisation that will be introduced in the definition of the fibre bundle in the following.
-Although a tangent space can be defined on any space defined as a manifold, here the vacuum is defined in the tangent space of $M$, as explained above.
+The vanishing curvature of the gauge field is understood if the gauge theory in the tangent space (of the spacetime) is also described by the tangent space of the principal bundle of the gauge.
+Thus, by treating both Lorentz and gauge symmetries equally, the local vacuum can be defined as having no curvature in either sense.
+Since a tangent space can be defined on any space defined as a manifold, here the vacuum is defined using these tangent spaces.
+What will be important in our calculations is that the connection is always zero at the contact point with the tangent space, where the local vacuum is defined.
 Above, we have described tangent vector bundles to introduce the concept of bundles.
@@
 [^6] Intuitively, the local direct product structure gives a point where one can start the discussion with a global symmetry (because they are given by the direct product) when describing a local symmetry.
+This procedure is one that we often see in field theory, whether or not there is a strict definition of local trivialization.
 When considering general fibre bundles, the direct product structure has to be introduced by hand.
@@
 Thanks to the introduction of a local direct product structure in the fibre, connections can be defined in a natural way.
-At this point, the introduction of a direct product structure may not seem to make sense as physics, but when one considers the role that local inertial systems actually play in general relativity, one has to consider that it does make sense as physics.
-Since in mathematics general relativity and gauge theories are described in a unified way, it seems to be natural to treat the two in a unified way.
+At this point, the introduction of a direct product structure may not seem to make sense as physics.
+However, when one considers the role that local inertial systems actually play in general relativity, one has to consider that it does make sense as physics.
+Since in mathematics, general relativity and gauge theories are described in a unified way, it seems to be natural to treat the two in a unified way.
+Here, the “gauge field with curvature” yields the Schwinger effect, while the “spacetime with curvature” yields the Unruh effect and Hawking radiation.
 We believe that essential differences are highlighted only when analyses are held on the same foundations as far as possible.
@@
 For cases where the mapping has internal degrees of freedom, a section is seen as one concrete form is selected.
-In physics, selecting the frame of a particular observer in special relativity is a section of the frame bundle.
+In physics, selecting the frame of a particular observer in special relativity gives a section of the frame bundle.
+For an accelerating observer, the observer’s frame moves on the frame bundle.
+This non-trivial section is called the moving frame.
+The range of applicability of the moving frame has been calculated in Ref.[12].
+It is only consistent in the range $\propto a^{-1}$, where $a$ is the accelerating rate of the observer.
+This point is very important, since the conventional calculation extrapolate well beyond this range and this may have led to an entanglement that should not have occurred.
+Hypotheses of this kind can only be tested by means of local calculations without extrapolation.
 It is particularly important not to confuse conventional “gauge fixing” with such a “section”.
@@
 We will later analyze particle production in the vicinity of the contact point of a tangent space, but be careful not to confuse “the section used to define the contact point of the tangent space” with “gauge fixing of quantum fluctuations”.
-Above, fibres have been introduced locally by means of local trivialization, and fibres on $U_{i}$ are laminated together to form a fibre bundle.
+Above, fibres have been introduced locally by means of local trivialization, and the fibres are laminated together to form a fibre bundle.
 When local trivialization varied from place to place, it was necessary to connect them by using the transformation.
@@
 For the same reason, the value of the connection is supposed to vanish at the contact point of the tangent space.
+The reason why non-trivial connections appear in the equations of an observer in accelerated motion in flat space is very easy to understand from a differential geometry point of view.
+This is because, although the space-time itself is flat, the observer’s equation sees a non-trivial section of the bundle.
 To explain the situation using a concrete example, we consider a two-dimensional real vector field $\phi$: $M\rightarrow\mathbb{R}^{2}$ as a section of a vector bundle $E$, where $M=\mathbb{R}^{3}$ and $F=\mathbb{R}^{2}$.
@@
 Although the transformation of the scalar field is trivial under the coordinate transformation, the effect of gravity is incorporated in a natural way as $d\phi$ is also a section of the frame bundle.
-The frame bundle is described below.
+The frame bundle will be described later.
 Given that Lorentz and gauge symmetries are treated as equivalence classes in field theory, one might think that the above discussion does not adequately describe the situation.
@@
 The simplest model uses the electromagnetic $U(1)$ gauge symmetry on a flat space-time and the curvature is introduced by a constant electric field.
-In this model, the manifold is static in the sense that the curvature is constant, but quantum theory expects a dynamical phenomenon on it.
+In this model, the manifold is static in the sense that the curvature is constant, but quantum theory expects a dynamical phenomenon (particle production) on it.
 The particle production in this model is called the Schwinger effect[3].
 In this case, the quickest way to avoid tedious discussions about manifolds is to use a powerful computational tool, the path integral[3, 13, 14].
-However, in this paper, we venture a primitive analysis based on field equations to look closely at what happens on the manifold.
+However, in this paper, we venture a primitive analysis based on field equations to look closely at what happens in differential geometry and manifolds.
 The analysis on the manifold using a scalar field is already given in Ref.[15, 16, 17].
-What is new in this paper is the analysis of the fermionic Schwinger effect as the Landau-Zener transition and its application to a time-dependent electric field.
-We show that the concept of the manifold is particularly important when considering a slowly varying electric field.
+The analysis of the fermionic Schwinger effect as the Landau-Zener transition and its application to a time-dependent electric field is new.
+Using this model and a slowly varying electric field, we show why the concept of the manifold and definition of local vacuum in the tangent space is particularly important.
 To understand the fermionic Schwinger effect as the Landau-Zener transition, we introduce the conventional decomposition of the Dirac fermion as[18, 19, 20]
@@
 Cosmological particle production after inflation has been discussed for time-dependent mass $m(t)$ in various situations.
-Such particle production is called preheating[18, 19].
+Such particle production is called preheating[21, 22], for which there is no serious need for careful discussions about locality.
+A serious discussion about locality is important when the system appears to be static and a symmetry is related to it.
+The most important topic in this section is how this relationship between locality and symmetry is described in differential geometry.
 It was first recognized in Ref.[23] that the Fermion preheating can be interpreted as the Landau-Zener transition[11], and the idea has been extended in Ref.[19, 20] to solve cosmological problems of particle production.
@@
 The reason for the somewhat roundabout approach here is that we want to examine the Stokes lines in the vicinity of the tangent space.
-To understand the structure of the Stokes lines, we use the Exact WKB (EWKB) developed in Refs.[21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
+To understand the structure of the Stokes lines, we use the exact WKB developed in Refs.[24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
 $\ddot{X}-\frac{\dot{\Delta}^{*}}{\Delta^{*}}\dot{X}+\left(-\frac{iD\dot{\Delta}^{*}}{\hbar\Delta^{*}}+\frac{i\dot{D}}{\hbar}+\frac{|\Delta|^{2}+D^{2}}{\hbar^{2}}\right)X=0.$ (3.27)
@@
 , the Stokes lines of the above equation coincide with the simple equation[32, 34]
-$\ddot{\hat{X}}$ $+$ $\frac{|\Delta|^{2}+D^{2}}{\hbar^{2}}\hat{X}=0.$ (3.33)
-When $\Delta(t)=\lambda t$ and $D(t)=D_{0}$, the above equation is a well-known problem of quantum mechanics (i.e, scattering by an inverted quadratic potential).
+$\ddot{\hat{X}}$ $+$ $\frac{|\Delta|^{2}+D^{2}}{\hbar^{2}}\hat{X}=0,$ (3.33)
+where $V(t)\equiv-(|\Delta|^{2}+D^{2})$ is called a “potential” of the “Schrodinger equation”.
+When $\Delta(t)=\lambda t$ and $D(t)=D_{0}$, the above equation is similar to a well-known problem of quantum mechanics (i.e, the scattering by an inverted quadratic potential).
+The solutions of $|\Delta(t)|^{2}+D(t)^{2}=0$ are called “turning points”.
+The main difference between the analysis using the exact WKB and the conventional analysis is that the local analysis in the vicinity of the Stokes line is mathematically well defined in the exact WKB.
+This distinguishing property of the exact WKB is the most important when considering local analysis.
 There are a few things to be considered with care when using this equation for solving the Schwinger effect.
-The first and the most important is the definition of the vacuum.
+The first and the most important is the definition of the local vacuum.
 In the usual analysis of a constant electric field, one will find a scattering problem by an quadratic potential, where the two vacuum states (in and out states at $t=-\infty$ and $t=+\infty$ where the electric field is supposed to disappear) are defined as asymptotic states.
@@
 However, as already explained, tangent spaces are naturally introduced in manifolds, and these tangent spaces have the property for defining the local vacuum states.
-Therefore, instead of losing the locality of the analysis by assuming the vacuum states at far away, we try to solve the problem by defining the vacuum on the tangent space.
-By definition, a feature of the tangent space is that the connection is zero in the space.
-This feature must be realized on the manifold at the point of contact.
-This is also a simple consequence of the local trivialization.
-This means that the gauge symmetry of each $U_{i}$ must be used carefully to make the connection vanish at the point.
-This makes the top of the quadratic potential coincide with the tangent space, as is illustrated in Fig.1.
+Therefore, instead of losing the locality of the analysis by assuming the vacuum states at far away, we try to solve the problem by defining the vacuum in the tangent space.
+As we have already mentioned, a feature of the tangent space is that the curvatur is zero in the space, and the connection must vanish at the point of contact.
+This is also a simple consequence of the local trivialization we have described above.
+This means that the gauge symmetry of each $U_{i}\in M$ must be used carefully to make the connection vanish at the point.
+Although it may not be trivial, one can see from the equation that this makes the top of the quadratic potential coincide with the tangent space in each $U_{i}$, as is illustrated in Fig.3.
 Figure 3: The typical “potential” of the equation is shown for $V(t)=1-t^{2}$.
@@
 The Stokes lines appearing on $U_{i}$ of $M$ are illustrated in the bottom picture.
+The top of the local “potential” must coincide with the contact point of the tangent space because of the local trivialization in each $U_{i}$.
+This makes the Stokes lines coincide at the contact point and makes each $U_{i}$ looks completely the same.
 Then, they are laminated using gauge transformations.
@@
 Similarly, since the tangent space is also a local inertial system, the velocity (or $k$) is assumed to be negligible to account for the inertial vacuum seen by the generated particle.
-[^10] In the Unruh effect, the vacuum is non-trivial when it is seen by the particle, not by someone else.
-Note that choosing a section of the observer (i.e, the inertial frame) does not kill the Lorentz symmetry of the theory.
-Similarly, the choice of the gauge in $U_{i}$ does not kill the gauge symmetry of the theory.
 This changes the definition of the electric field $E_{0}$ in the equation since the original $E_{0}$ was defined for an observer in the laboratory.
@@
 Now we have defined everything in the vicinity of the contact point of the tangent space.
-Note that the “decoupling” of the equations performed above cannot be defined at the point of contact with the tangent space, as the non-diagonal element disappears from the matrix and the equations are already decoupled there.
+Note that the “decoupling” of the equations performed above cannot be defined at the point of contact with the tangent space, as the non-diagonal elements disappear from the matrix and the equations are already decoupled there.
 This shows that there is no mixing in the tangent space by definition.
 As discussed in the definition of manifolds, the Stokes phenomenon should be considered on an open set $U_{i}$ defined in a neighborhood of $p\in M$, and it is the physics around $p$ that is relevant for mixing solutions.
-Therefore, the Stokes phenomenon is considered here in a neighborhood where the point $\Delta=0$ is slightly avoided.
+Therefore, the Stokes phenomenon is considered here in a neighborhood.
 The above equations show that the real-time axis traverses the Stokes line in the vicinity of the tangent space on every $U_{i}$[16].
@@
 This is the same situation as the asymptotic states.
-In the same way as the correspondence with the vacuum is considered for the asymptotic states, the correspondence with the vacuum solution here is considered in the vicinity of the tangent space.
+In the same way as the correspondence with the vacuum is considered for the asymptotic states, the correspondence with the vacuum solutions here is considered in the vicinity of the tangent space.
 Here, one might notice that the frame prepared for the generated particles of the Schwinger effect actually represents an accelerated frame called the moving frame.
@@
 This topic will be discussed in the next section.
-In the remaining part of this section, we are going to look at the time-dependent case in a little more detail.
+In the remaining part of this section, we are going to look at the time-dependent case in a little more detail to show a simple example in which definition of the local vacuum is crucial for the calculation.
 3.2 Time-dependent $E(t)$
-We will now try to show an example where it makes no sense at all as physics to find a naive global solution to an equation defined on an open set $U_{i}$.
-It naturally depend on the situations whether a meaningful result can be obtained by such solutions when extrapolating the equations originally defined on $U_{i}$.
-However, almost all papers do not mention this point at all.
+We will now try to show an example where it makes no sense at all as physics to find a naive global solution by considering extrapolation to an equation defined on an open set $U_{i}$.
+It naturally depend on the situations whether a meaningful result can be obtained by such extrapolation.
+However, as far as we know, there has been no paper in which this point is mentioned.
 In order to understand the problem without ambiguity, consider the case where the electric field changes gradually and there is no significant back-reaction from particle production.
 If the electric field oscillates, our assumption here is that the time period of the oscillation is much longer than the width of $U_{i}$.
-[^11] If the period of the oscillation is short, averaging over a period should be considered for the observation.
-In addition, the Stokes phenomenon in mathematics occurs on the Stokes “line” on the complex plane, but in physics, it will be better to assume an intrinsic width of the line.
-Therefore, the width of on an open set $U_{i}$ is assumed to be wider than the intrinsic width of the Stokes line.
-Replacing $E_{z}(t)=E_{0}$ with $E_{z}(t)=\alpha t$, the problem becomes scattering by a “quartic” potential.
+Replacing the constant electric field $E_{z}(t)=E_{0}$ with a slowly varying electric field $E_{z}(t)=\alpha t$, the problem becomes scattering by a “quartic” potential $V(t)\propto-t^{4}$.
 The solution to this problem and the Stokes lines have been studied in great detail by Voros[25], and it has been found that the Stokes lines act away from the origin.
@@
 Details concerning conventional particle production can be found in Ref.[23].
-Unlike the case of the constant electric field, the gauge ambiguity cannot naively shift the position of the quartic potential.
-Also, although intuitively the probability of particle production should change gradually if the electric field changes slowly, such calculation (scattering by the quartic potential solved for asymptotic states) will not change in such a way.
-Although it is immediately obvious to anyone who looks at the equation, we point out that the solutions of the equation, if they are used for the asymptotic states, can not solve the problem at all.
-It is easy to say that the setting of the problem is bad, but here we are going to try to clarify our understanding a little bit more.
-For our purpuses, it is important to go back to the definition of the manifold and solve the problem using local analysis.
+What is important here is that unlike the case of the constant electric field, the gauge ambiguity cannot naively shift the position of the quartic potential.
+Also, although intuitively the probability of particle production should change gradually if the electric field changes slowly, a conventional calculation (scattering by the quartic potential solved for asymptotic states) will not change in such a way.
+Although it shoule be immediately obvious to anyone who looks at the equation, we point out that the solutions of the equation, if they are used for the asymptotic states, can not solve the problem at all, even if the solutions are exact.
+It is easy to finish the discussion by saying that the setting of the problem is bad, but here we are going to try to clarify our understanding a little bit more.
+For our purpuses, it is important to go back to the definition of the manifold and solve the problem in local.
 Figure 4: The typical Stokes lines of the scattering problem by a quartic potential are shown for $V(t)=1-t^{4}$. In contrast to the quadratic potential, the Stokes lines are not crossing the origin.
-Let us now reconsider the above question along the lines of the manifold and the differential geometry.
+Let us now rephrase the above question along the lines of the manifold and the differential geometry.
 First, consider an open set $U_{i}$ in the neighborhood of $t=t_{p}$ and let $A_{z}(t)=\frac{1}{2}\alpha t^{2}$ be expanded at $t=t_{p}$.
 Then, we have the “local potential” defined for $U_{i}$ as
-$V(t)|_{U_{i}}$ $\simeq$ $e^{2}\hat{E}_{z}(t_{p})^{2}(t-t_{p})^{2},$ (3.34)
-where by using the gauge of $U_{i}$ the connection is set to zero at the contact point $t=t_{p}$.
-This equation clearly meets the above requirement.
+$V(t)|_{U_{i}}$ $\simeq$ $m^{2}-e^{2}\hat{E}_{z}(t_{p})^{2}(t-t_{p})^{2},$ (3.34)
+where by using the gauge of $U_{i}$ the connection ($A_{\mu}$) is set to zero at the contact point $t=t_{p}$.
+$k\simeq 0$ is also expected for the inertial frame of the particle.
+Here the tangent space and the local vacuum is defined at $t=t_{p}$.
+This equation clearly meets the above requirement if the local vacuum is defined in the tangent space.
+See also Fig.5.
+The particle production can be calculated locally on each $U_{i}\in M$ by using $V(t)|_{U_{i}}$ and the production rate gradually changes with time.
 In this way, the generation rate can be calculated for the local time ($t_{p}$) on the local open set ($U_{i}$).
-There could be critical opinions about not defining asymptotic states for solving field equations to analyze particle production, but at least for cases like the one discussed here, the use of the tangent space seems to be inevitable.
-We would like to stress the importance of using two different definitions of the vacuum for different purposes.
-Furthermore, if the electric field changes much more rapidly and is treated as quantized, then the above treatment as a “background field” would no longer be appropriate and the field equations of the gauge field come into play.
-Then, we cannot imagine a better alternative to Feynman diagrams and path integrals.
-As we have already seen in the simple example of a slowly changing electric field, it is very important and sometimes quite essential to solve the field equations locally on the manifold.
-Of course, when determining the averaged particle production rate in the case of repeated rapid oscillations or when the electric field is instantaneous, local evaluation of the particle production rate is useless for the experimental observation.
-Then, to get the required results, one should consider asymptotic states and averaging over a period.
+For cases like the one discussed here, the use of the tangent space to define the local vacuum is inevitable.
+The conventional calculation defining the asymptotic vacuum at a distance could be convenient but not rigorous in mathematics.
+We would like to stress the importance of understanding benetits and risks of the definitions of the vacuum.
+As far as we can understand, the serious risk of defining the vacuum in the asymptotic state has not been explained in detail so far for the Schwinger effect.
+Figure 5: The upper picture shows the Stokes lines of the solution when the equation defined at $t=0$ is extrapolated to infinity.
+The lower picture shows the Stokes lines when equations are defined for each $U_{i}$ using local trivialization.
+One can see that the production rate is gradually changing with time (as the distance between two turning points are changing with time).
+As we have already seen in the simple example of a slowly changing electric field, it is very important and sometimes quite essential to solve the field equations locally on the manifold by defining a local vacuum in the tangent space.
+Of course, when determining the averaged particle production rate in the case of rapid oscillations or when the electric field is sharply instantaneous, local evaluation of the particle production rate should be useless for experimental observation.
+If the electric field changes rapidly, the field equation becomes a simultaneous differential equation with the electromagnetic field equation.
+The Stokes phenomena in such cases can be extremely complex and require appropriate approximations.
 [^13] Analysis of these topics using the Stokes lines can be found in Ref.[36].
 See also Refs.[37, 38]..
-What we have highlighted in this section is the case in which local calculation is needed to obtain correct results while a “naive” application of the conventional asymptotic states gives clearly a wrong answer.
-This distinction has never been explicitly recognized.
-We believe that the importance of the option of defining the vacuum in the tangent space has been made very clear by the simple model described in this section.
+What we have highlighted in this section is the case in which local calculation is crucial to obtain correct results while a “naive” extrapolation of the equation to the outside of the defined area gives clearly a wrong answer.
+This distinction has never been explicitly recognized so far.
+We believe that the importance of the option of defining the local vacuum in the tangent space has been made very clear by the simple model described in this section.
 Let us now look at how such a definition of the vacuum has implications in relativity.
+As mentioned above, particles created by the Schwinger effect can also be affected by the Unruh effect.
+In this case, the question is whether the two effects are independent phenomena or whether we are simply seeing the same thing.
+In the following, we present the solution to this problem using differential geometry and the exact WKB.
 4 The Unruh effect and Hawking radiation on manifolds
-The previous section dealt with the case where there is no qualitative ambiguity in defining the local vacuum, but in general relativity, even after the connection and the tangent spaces are determined, there are still further degrees of freedom left in the vierbein, which leaves ambiguity in defining the local vacuum.
+The previous section dealt with the case where there is no qualitative ambiguity in defining the local vacuum, but in general relativity, even after the connection (metric) is determined, there are still further degrees of freedom left in the vierbein, which leaves ambiguity in defining the local vacuum.
 The most obvious difference is between the Lorenz frame and the local inertial frame.
@@
 However, they are distinguished by the vierbein.
-The different choices of the “vacuum” is essential in the search for the Stokes phenomenon[13].
+The difference in the vierbein is essential in the search for the Stokes phenomenon of the Unruh effect[16].
 We describe the frame bundle further below.
 Associated with a tangent bundle $TM$ over $M$ is a principal bundle called the frame bundle $LM\equiv\bigcup_{p\in M}L_{p}M$ where $L_{p}M$ is the set of frames at $p$.
-The bundle $T_{p}M$ has a natural coordinate basis $\{\partial/\partial x^{\mu}\}$ on $U_{i}$ and a “frame” $u=\{X_{1},...,X_{m}\}$ at $p$ is expressed by the non-coordinate basis
+Here we have a natural coordinate basis $\{\partial/\partial x^{\mu}\}$ and a “frame” $u=\{X_{1},...,X_{m}\}$ at $p$ is expressed by the non-coordinate basis
 $X_{\alpha}$ $=$ $X^{\mu}_{\alpha}\left.\frac{\partial}{\partial x^{\mu}}\right|_{p}$ (4.5)
@@
 The following point is very important for our discussion.
-A natural coordinate basis is prepared on the surface of $M$ and the inertial system is defined using a non-coordinate basis.
+A natural coordinate basis is prepared for $U_{i}\in M$ and the inertial system is defined using a non-coordinate basis in the tangent space.
 This procedure naturally gives a notion of the “moving frame”[12], as the inertial frame seems to be rotated from time to time by the Lorentz transitions of the vielbeins, somewhat like spinning tea cups in amusement parks.
-Thomas precession is explained by the fact that multiple Lorentz transformations with different directions produce a rotation of the intrinsic space of the observer.
+[^15] One can see more explanations in Ref.[12], in which figures for the moving frame can be found.
+The Thomas precession is explained by the fact that multiple Lorentz transformations with different directions produce a rotation of the intrinsic space of the observer.
 If the observer stays in the same frame (no acceleration), there is no motion in the direction of the fibre of the frame bundle.
-In this case, the connection vanishes by definition because there is no transformation when laminating.
+In this case, the connection vanishes by definition.
 Therefore, for an inertial observer, the distinction between coordinate and non-coordinate systems will be quite ambiguous.
-However, if one wants to describe an observer in accelerated motion on flat space-time, one has to define the local inertial frame for the observer, which moves on the frame bundle in a non-trivial way.
+However, if one wants to describe an observer in accelerated motion on flat space-time, one has to define the local and subjective inertial frame for the observer in the tangent space, which moves on the frame bundle in a non-trivial way.
 This defines the section of the observer on the frame bundle.
 This means that on this section, one has to laminate the bundle by using non-trivial coordinate transformations.
+The Thomas precession is explained that the lamination by the Lorentz transformation causes rotation of the intrinsic space.
 This (the observer’s non-trivial section of the frame bundle) introduces the connection for the observer, although the space-time is flat.
@@
 $(e_{A})_{\alpha}^{\mu}$ $=$ $\left(\begin{array}[]{cc}\cosh a(\tau-\tau_{A})&\sinh a(\tau-\tau_{A})\\ \sinh a(\tau-\tau_{A})&\cosh a(\tau-\tau_{A})\end{array}\right).$ (4.8)
-Indeed, for such constant acceleration, the vielbeins have completely the same form for any time $(\tau)$.
+Indeed, for such constant acceleration, the vielbeins have completely the same form for any time $(\tau)$. At $\tau=\tau_{B}$, we have
+$(e_{B})_{\alpha}^{\mu}$ $=$ $\left(\begin{array}[]{cc}\cosh a(\tau-\tau_{B})&\sinh a(\tau-\tau_{B})\\ \sinh a(\tau-\tau_{B})&\cosh a(\tau-\tau_{B})\end{array}\right).$ (4.11)
 The transformation $(e_{A})_{\alpha}^{\mu}\rightarrow(e_{B})_{\alpha}^{\mu}$ on the frame bundle is the Lorentz transformation;
@@
 Note that the situation is very similar to the Schwinger effect for a constant electric field.
-In the Schwinger effect, the potential has been shifted by a gauge transformation; in the Unruh effect, the Lorentz transformation gives vielbeins of exactly the same shape.
+In the Schwinger effect, the potential has been shifted by gauge transformation and the equation looks always the same on each $U_{i}$.
+In the Unruh effect, the Lorentz transformation gives vielbeins of exactly the same shape on each $U_{i}$.
 In both cases, the observer is always looking at the same physics on a static manifold.
-Of course, the mathematical description of the frame bundle does not require an observer, but in physics an observer is inevitable.
-In this case, the observer is nothing but a section of the frame bundle.
-Previously, we have introduced $d\phi$ to describe covariant derivatives of the gauge symmetry.
-We have seen that both $\phi$ and $d\phi$ are the sections of the vector bundle, and the connection is introduced to make $d\phi$ covariant by the gauge symmetry.
-Now we can see that $d\phi$ is also a section of the frame bundle.
-If an observer stays in the same frame, there is no (non-trivial) connection required for $d\phi$.
-However, when an observer is accelerating, $d\phi$ traverses the frame bundle in the direction of the fibre and its motion causes Lorentz transformation.
-This is called “the moving frame”.
-Is local analysis possible also for the Unruh effect, if we follow the previous calculations of the Schwinger effect?
-Our answer is “No”.
-What is important here is that the definition of the covariant derivative uses the Lorenz frame in which the vierbein is diagonalized in the neighborhood.
+Our primary question here is “Is the same local analysis possible also for the Unruh effect, if we follow the previous calculations of the Schwinger effect?” Our answer is “No”.
+What is important here is that the mathematical definition of the covariant derivative uses the Lorenz frame in which the vierbein is diagonalized in the neighborhood.
 On the other hand, as is shown explicitly above, the vierbein of the local inertial frame is diagonalized only at the point.
-(Note that the off-diagonal elements vanish since $\sinh a(\tau-\tau_{A})=0$ at $\tau=\tau_{A}$.) Since the covariant derivatives are defined for the Lorentz frame while the Unruh effect is defined for the inertial frame, it seems impossible to examine the Stokes phenomena of the Unruh effect directly (and locally) in terms of the field equations[13].
+(Note that the off-diagonal elements vanish at the point since $\sinh a(\tau-\tau_{A})=0$ at $\tau=\tau_{A}$.) Since the covariant derivatives are defined for the Lorentz frame while the Unruh effect is defined for the inertial frame, it seems impossible to examine the Stokes phenomena of the Unruh effect directly (and locally) in terms of the field equations[16].
 This mismatch has prevented the local analysis of the Unruh effect for a long time.
-Now consider the physics that observers see in the Unruh effect.
+Now we consider the physics that observers see in the Unruh effect.
 As far as the acceleration seen by the observer is constant, the physics seen by the observer is indistinguishable at any time due to the equivalence classes defined for the manifold.
@@
 If a dynamical effect (particle production) is manifested in such a situation, it must be explained by the vierbein of the local inertial frame.
-Extrapolating the coordinates to infinity and considering a global map as the Bogoliubov transformation is conceptually unacceptable, even if it could be plausible as a method[3].
+Extrapolating the coordinates to infinity and considering a global map as the Bogoliubov transformation is conceptually unacceptable, even if it could be commonly used as a method[6].
 In fact, it is known that such methods can lead to unnatural entanglements appearing between regions that should be uncorrelated.
@@
 Rindler coordinates are a specific set of coordinates used in the context of special relativity to describe the motion of an observer undergoing constant proper acceleration in flat spacetime.
-Thus, the Rindler coordinates form a coordinate chart that covers a specific region of Minkowski spacetime known as the Rindler wedge.
+Thus, the Rindler coordinates form a coordinate chart (not a frame) that covers a specific region of Minkowski spacetime known as the Rindler wedge.
 Following the concept of manifolds, the simplest local inertial system is defined as a tangent space in the neighborhood of the point with zero velocity in Rindler’s coordinates.
-We start with the conventional Rindler metric[3, 9, 38];
+If one starts with the conventional Rindler metric[6, 12, 41]; [^16] Note that we do not use the metric in our calculations.
+The reason for showing a metric here is just to explain that the Stokes phenomenon cannot be calculated by the metric.
 $ds^{2}$ $=$ $-\left(1+ax\right)^{2}dt^{2}+dx^{2},$ (4.15)
-where $x=a^{-1},t=0$ is the place for defining the simplest local inertial space.
-If one tries to define a local inertial system for any other point, one has to consider a distorted coordinate system due to Lorentz transformation.
-[^14] One can see more explanations in Ref.[9], in which figures for the moving frame can be found.
-As is immediately apparent from the metric, the field equation does not yield the local Stokes phenomenon as was the case with the Schwinger effect[12].
+one will not be able to find the required Stokes lines.
+As is immediately apparent from the metric (and the equations), the field equation does not yield the local Stokes phenomenon as was the case with the Schwinger effect[15].
 This indicates that there may be a fundamental error in the way of setting the problem.
-So far, many researchers have truncated locality here and moved on to find a solution in a global space[2, 3], but here we shall stick to the locality of the issue.
+Here, we shall stick to the locality of the issue.
 Since the metric (covariant derivatives) does not explain the local Stokes phenomenon, the only way left for us is to use the vierbein.
-The problem is that even if the vierbein is used for the calculation, the same trivial result is obtained once the field equations are written down.
+The problem is that even if the vierbein is used for the calculation, the same trivial result is obtained once the field equations are written down, as the Lorentz frame and the inertial frame are not distinguishable by the metric.
 Therefore, the vierbein must be used without going through the field equations.
@@
 The vacuum is observed here by the particle itself, which is ejected from “the vacuum”.
+[^17] Alternatively, one can introduce the Unruh-DeWitt detector as an observer.
+The problem has been solved in Ref.[17] by paying serious attention to the locality and the differential geometry.
+The result is consistent with this paper.
 Since such particles have no momentum in their unique frame, we can neglect the $x$-dependent component of the vacuum solution.
@@
 […]
-which have the required Stokes phenomenon.
+where $dt$ is needed for the calculation, even if $\omega$ is a constant. The situation is illustrated in Fig.6.
+Figure 6: The local vacuum defined in the tangent space is seen by an accelerating observer using the vierbein.
+The Stokes phenomenon on the Moving frame is illustrated in the bottom.
+The situation is always the same for the observer if the acceleration is constant.
+The observer sees always the same Stokes phenomenon on the moving frame, where the subjective vacuum for the observer is defined in the tangent space using the inertial frame of the observer.
+We have already analyzed the Stokes phenomenon of the above function in detail in Ref.[15, 17].
+As the details of the Stokes phenomenon of the solution are not the issue of this paper, we will present only the results here.
 See Ref.[15, 17] if the reader is interested in the details about the Stokes phenomenon.
 One thing that should be noted is that the result does not meet the conventional (global) calculation by a factor of two.
-This point has to be explained.
+This point has to be explained here.
 To make the point clear, we consider the entanglements that appear in conventional calculations[6].
-In our calculation, a difference of the factor of 2 appears in the non-perturbative factors emerging from the Stokes phenomenon.
-With entanglement, the probability ($P_{1}$ for a particle production) is squared because there are two particles produced.
+In the conventional calculation, the factor of 2 is added when taking a trace on the entanglements in the distant wedge.
+This is because the particle creation in the right wedge is always accompanied with that in the left wedge as $e^{-\pi\omega/a}b_{\bm{\mathrm{k}}}^{L\dagger}b_{\bm{\mathrm{k}}}^{R\dagger}|0_{L}\rangle\otimes|0_{R}\rangle$, where $b_{\bm{\mathrm{k}}}^{L\dagger}$ and $b_{\bm{\mathrm{k}}}^{R\dagger}$ are the creation operators in the left and the right wedges, respectively.
+One can see that there is a duplication of the factor, which can be separated as $\left(e^{-\pi\omega/2a}b_{\bm{\mathrm{k}}}^{L\dagger}\right)\left(e^{-\pi\omega/2a}b_{\bm{\mathrm{k}}}^{R\dagger}\right)|0_{L}\rangle\otimes|0_{R}\rangle$.
+This duplication does not appear if the calculation is local.
+The entanglement appeared because of the extrapolation beyond the applicable range of the moving frame that defines the observer’s vacuum.
+Also, to be more rigorous, defining a vacuum in the chart is not a preferred method.
+In our calculation, we have the factor $e^{-\pi\omega/2a}$ because it is derived from the local Stokes phenomenon.
+For a local pair creation, we will find $e^{-\pi\omega/a}$ but in our case both particles are observed and there is no trace-out.
+If we assume the entanglement between the distant wedge, our result can reproduce the usual Unruh effect calculations, but since our local calculation is rigorous as mathematics, there is no need to assume such entanglement with a distant wedge.
+This is the Bogoliubov coefficient of the state mixing in the vacuum when the vacuum is seen by an accelerating observer.
+Let us be more precise.
+Noting that the vierbein connects the inertial frame (vacuum) and the observer, we rewrite $\phi$ in the vacuum as
+[…]
+where $\omega$ is constant but $dt$ is required to introduce the vierbein in an explicit form.
+If a vierbein gives $dt=(e(\tau))^{t}_{\tau}d\tau$, mixing of the solutions (when it is seen by an accelerating observer whose proper time is $\tau$) after crossing the Stokes line is written by
+[…]
+Then, we have
+[…]
+[…]
+which suggests that the particle production is seen by $\beta_{\pm}\neq 0$.
+We call this factor the Bogoliubov coefficient, and this factor should be considered as the Boltzmann factor of the Unruh effect.
+With entanglement, the probability ($P_{1}$ for a particle production in our calculation) is squared because there are two particles produced but only one particle is observed.
 This gives $P_{entangled}=P_{1}^{2}$ in our calculation.
-In this way, our calculations will give the same results as the conventional calculations if the entanglement is included by hand.
-In short, if we “assume” entanglement, the two calculations coincide.
-So what happens if we use our calculations for Hawking radiation?
-Conventional calculations of the Unruh effect treat infinite inertial space as real, so the existence of such spaces on curved manifolds cannot be assumed.
-On the other hand, our calculations are based on the definition of manifolds, so the same calculations can be performed in the vicinity of the black hole horizon.
+In this way, our calculations will give the same result as the conventional calculation if the entanglement is included by hand.
+In short, if we “assume” entanglement, our calculation coincides with the conventional calculation, but obviously our local calculation does not require the entanglement.
+In our previous paper[17], we have calculated the Boltzmann factor of the Unruh-DeWitt detector and found the same result.
+Therefore, our local calculations are consistent between the Unruh effect and the Unruh-DeWitt detector, but are not consistent with conventional calculations.
+The discrepancy is explained by the entanglement.
+Our local analyses in terms of the differential geometry and the exact WKB allow calculating the probability of single particle observation as the single particle production.
+Then, our results differ by a factor of two from the usual Unruh Effect and the Unruh-DeWitt detector calculations.
+So what happens if we apply our calculations to Hawking radiation?
+Conventional calculations of the Unruh effect treat infinite inertial space as real, so the existence of such spaces cannot be assumed for Hawking radiation.
+On the other hand, our calculations are always locally defined, so the same calculation of the Unruh effect can be performed in the vicinity of the black hole horizon.
 When a pair of particles is created in the vicinity of the horizon, one particle inside the horizon can have negative energy (when it is seen from the outside), and the other in the outside can have positive energy.
-Again, as there are two particles produced, the total production probability is squared to obtain results consistent with Hawking radiation[12, 13].
-Thus, if all calculations are performed faithfully to the definition of the manifold, strong doubts will arise about the entanglement of the Unruh effect.
-[^15] We are not claiming that our analysis is able to provide a proof that there is no entanglement in the global calculations.
-It is an indisputable fact that entanglement appears in the global calculation.
-In the following, we will consider the Unruh effect when the acceleration changes, which is similar to the slowly changing electric field in the Schwinger effect.
-We are aware of papers claiming that the Unruh effect can be solved for various accelerating motions of the Unruh-DeWitt detector by adjusting the classical orbit of the detector.
-Again, our analysis cannot provide a proof that these results are wrong, but we believe we are giving a better alternative.
-Simply because the general calculations of the Unruh-DeWitt detector are somewhat different from the Unruh effect itself, we have treated them as different.
-See Ref.[14] for more details about what the local calculations of the Unruh-DeWitt detector look like if it is defined on a manifold, where we found the same factor of two discrepancy between the conventional global analysis and our local analysis.
+Again, as there are two particles produced but only one particle is observed as radiation, the total production probability for a particle radiation is squared to obtain results consistent with Hawking radiation[15, 16].
+In this case, unlike the Unruh effect, we do not have to assume unnatural entanglement at a distance.
+Thus, if all calculations are performed faithfully to mathematics of the differential geometry, strong doubts will arise about the entanglement of the Unruh effect.
+[^18] We are not claiming that our analysis is able to provide a proof that there is no entanglement in the conventional global calculations.
+It is an indisputable fact that entanglement appears in the global calculation, although the mathematical consistency is not reliable because of the extrapolation.
+For completeness, we will now consider the Unruh effect as the acceleration changes, which is similar to the slowly changing electric field in the Schwinger effect.
+Simply because the general calculations of the Unruh-DeWitt detector are somewhat different from the Unruh effect itself, we will treat them as different.
+See Ref.[17] for more details about what the local calculations of the Unruh-DeWitt detector look like if it is defined locally using the differential geometry and the exact WKB.
 Note also that the Unruh-DeWitt detector requires an explicit interaction with the detector to be included, which is unlikely to be applicable directly to Hawking radiation.
 4.1 When the acceleration rate is slowly time-dependent
-For later calculations, let us first derive the Rindler coordinate for the case of constant acceleration using a somewhat roundabout approach.
+Let us first derive the Rindler coordinate for the case of constant acceleration using a somewhat roundabout approach.
 For simplicity, we restrict the motion to the x-axis direction.
@@
 Since the function $T$ of $T^{\prime}$ (i.e, $T(T^{\prime})$) is periodic in the direction of the imaginary axis of $T^{\prime}$, we can expect any vacuum function $F(T)$ described by the observer’s time $T^{\prime}$ to be periodic in the observer’s complex time.
-Intuitively, this suggests that what the accelerating observer sees in the vacuum is expected to be thermal.
 Here a simple question would arise.
@@
 Obviously, if we consider $\beta\neq 0$, the periodicity of the function $T(T^{\prime})$ is not a simple imaginary.
-Such “coordinate systems” do not have the special properties of the Rindler coordinates.
-Therefore, Unlike the conventional Unruh effect for a constant acceleration, the vierbeins cannot be moved to the same form by the Lorentz transformation.
+The above “coordinate system” for the slowly varying acceleration do not have the special properties of the Rindler coordinates.
+More specifically, the situation cannot be always the same for the observer.
+This means that unlike the conventional Unruh effect for a constant acceleration, the vierbeins cannot be moved to the same form by the Lorentz transformation.
 This situation is quite similar to the case which appeared when we have considered the weakly time-dependent electric field in the Schwinger effect.
@@
 In the present case we have $m<0$, while for an pendulum it becomes $m>0$.
-We comment on the case of solving it at other times.
+We comment on the case of solving it at other times ($t\neq 0$).
 If we consider the Unruh effect at $t=t_{0}$, the inertial condition is now $v(t_{0})=0$.
@@
 The two phenomena, which at first sight appear to be the same, turn out to be caused by completely different sources when one looks at the local structure of the manifold.
-As we expect negative reactions to the idea of defining the vacuum in the tangent space, this part of the article was explained in particular detail.
-In this paper, the case of a gradual change in curvature has been considered as an example where the correct answer can only be obtained when the vacuum is defined on the tangent space.
-In our local analysis, the entanglement of the Unruh effect appears to be an apparent one due to the extrapolation of the coordinate system, but this will require further multifaceted verification.
-The structure of the elliptic function for the complex time, manifested in the Unruh effect, is commonly found in curved spacetime.
-Further analysis is needed to determine what implication can be gained when such a structure is incorporated into a manifold.
-In physics, it is well known that the connection of the electromagnetic $U(1)$ and the metric can be literally unified into the Kaluza-Klein theory.
-Although in our discussion of the Unruh effect it might have seemed that the Stokes phenomenon was not able to be derived from the metric, it is clear from the Kaluza-Klein theory that it can appear from the metric once they are actually embedded in the Kaluza-Klein theory.
-On the other hand, concerning the Unruh effect, the situation remains the same in Kaluza-Klein theory, as the Unruh effect is still hidden in analyses using covariant derivatives.
-The underlying cause of the Unruh effect is that the inertial system is attached in a twisted manner, which cannot be detected by the covariant derivatives (at least locally).
-We hope that the local analyses presented in this paper will help people understand the physics of the quantum field theory on curved manifolds.
+As the idea of defining the vacuum in the local tangent space could be unfamiliar in the analyses of the Schwinger effect and the Unruh effect, this part of the article was explained in particular detail.
+In this paper, the case of a gradual change in curvature has been considered as an example where the correct answer can only be obtained when the vacuum is defined in the local tangent space.
+In our local analysis, the entanglement of the Unruh effect appears to be an apparent one due to the extrapolation of the coordinate system.
+On the other hand, our result agrees with Hawking radiation, where the basic procedures of differential geometry can be used.
+We will summarise the reasons why the factor of 2 difference appeared between the conventional calculation of the Unruh effect and ours.
+As we have described in detail in Sec.4, the direct reason for this difference is the entanglement between distant wedges.
+This entanglement does not actually exist.
+The problem arose because of the following problems of the conventional calculation.
+  • The moving frame was not moving. This ruins the mathematical consistency of the calculation.
+  • The Rindler “vacuum” was defined in the chart, but as we have described in detail in Sec.2, the chart is not a good place for defining a vacuum.
+One must remember how Lie algebra is introduced in differential geometry.
+  • The Rindler chart describes eternal acceleration, but the observer cannot be accelerated eternally in a flat spacetime.
+Mathematically, there is no problem at this point, as the chart can be chosen in any way.
+The problem arises when one ignores mathematical consistency and behaves as if the chart is “real”.
+  • It must be remembered that the “chart and frame” cannot be used outside the application area.
+When dealing with an accelerating observer, either frame or chart has to be Lorentz-transformed.
+Normally the frame is Lorentz-transformed and the frame is called the moving frame.
+Otherwise the mathematical consistency will be lost outside the application area[12].
+  • The most serious problem with defining a vacuum in the chart appears when information about the far future or the far distance (arbitrarily defined by the people who have chosen the chart) is used for the calculation.
+  • In the conventional calculation, the quantum states are assumed to be coherent over the whole space-time on the Rindler chart.
+The information in the far distance is used because of the coherence.
+The entanglement of the Unruh effect appears in this way.
+Of course, under normal circumstances, it would be inconceivable to define a state that is coherent to infinity on a chart that is merely a landmark.
+  • If the calculation is faithful to differential geometry and no “assumption” is made in physics, the information to be used in physics is normally in the frame and the vierbeins, not in the chart.
+Despite its critical importance in this field, differential geometry has not been used in good faith in the analysis of the Unruh effect.
+The reason is that the procedures of the conventional calculation cannot agree with the basic concepts of differential geometry.
+If one attempts to solve the problem faithfully to the concepts of differential geometry, one will find that tools are needed to derive the Unruh effect locally.
+The fact that such tools were not discovered has postponed the issue of the Unruh effect.
+The discovery of the local analysis by the exact WKB is a breakthrough in this sense.
+Now we would like to rephrase our most important conclusion. “The entanglement between distant wedges of the Unruh effect is an illusion caused by a computational glitch.”
+We hope that our local analyses presented in this paper will help people understand more clearly the physics of the quantum field theory on curved manifolds.